### v1

# Orthogonalized Policy Optimization Decoupling Sampling Geometry from Optimization Geometry in RLHF

Abstract
Recent alignment methods for large language models, including PPO, DPO, and IPO, are often presented as distinct algorithms. In this work, we show that many of these approaches implicitly conflate two fundamental and independent design choices: (i) the sampling geometry, which determines which samples dominate the gradient signal, and (ii) the optimization geometry, which determines how deviations in value are penalized.
We formalize this observation by expressing alignment as the minimization of a generalized distance between policy energy and target energy, parameterized by an $\alpha$-divergence-based sampling weight and a Bregman divergence-based value metric. We demonstrate that the commonly used KL divergence induces an exponential penalty on unbounded value signals, leading to numerical instability and vanishing gradients in high-confidence regimes.
To address this issue, we propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples sampling geometry from optimization geometry. By combining $\alpha$-weighted importance sampling with a $\chi^{2}$-induced quadratic regularization in ratio coordinates, OPO yields a simple and well-conditioned objective with linear gradient dynamics. This formulation maintains stable optimization while preserving peak-seeking behavior, and avoids gradient saturation even when model confidence is high. Our analysis positions OPO as a unifying perspective on existing alignment methods and provides a principled foundation for robust reasoning-oriented training.
1 Introduction
The field of Large Language Model (LLM) alignment has seen a proliferation of objective functions—PPO [2], DPO [1], IPO [7], GRPO [4]—each demonstrating effectiveness in specific settings. While these methods appear distinct, we observe that they share a common underlying structure that can be decomposed along two independent axes.
Observation: Implicit Conflation of Two Design Choices.
We identify two fundamental degrees of freedom in alignment objectives:
  1. 1.
    Sampling Geometry: Which samples should dominate the gradient? This determines whether optimization is "average-case" (covering the target distribution) or "peak-seeking" (concentrating on high-reward modes).
  2. 2.
    Optimization Geometry: How should deviations in value be penalized? This determines the functional form of the gradient with respect to the policy’s value signal.
In existing methods, these two choices are often coupled through a single divergence measure (typically KL). This implicit conflation can lead to undesirable interactions: for instance, adjusting the KL coefficient to control exploration simultaneously alters the curvature of the optimization landscape.
Diagnosis: The KL-Induced Exponential Penalty.
The standard KL-regularized RL objective induces an exponential penalty on value signals via the log-sum-exp conjugate. While this is well-suited for probability distributions, the Advantage signal in RLHF is an unbounded real value. We show that this mismatch leads to:
  • Numerical Instability: The exponential term can grow unboundedly.
  • Gradient Saturation: In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
Contribution: Orthogonalized Policy Optimization (OPO).
We propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples the two design axes. The key insight is to work in ratio coordinates rather than log-ratio coordinates, where the Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty. The OPO objective is:
$\mathcal{L}_{\text{OPO}}=-\sum_{y\in S}\omega_{\alpha}(y)\cdot v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2},\quad v_{\theta}(y)=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
where $\omega_{\alpha}(y)$ is an $\alpha$-divergence-derived importance weight controlling sampling geometry, and $\frac{\mu}{2}v^{2}$ is the Pearson $\chi^{2}$-induced quadratic regularization controlling optimization geometry. Intuitively, $\omega_{\alpha}$ increases the influence of rare high-quality samples while retaining a controllable interpolation to uniform weighting.
This decoupling allows independent control over "what to learn" (via $\alpha$) and "how to penalize" (via $\mu$), avoiding the entangled dynamics of KL-based methods. We show that this formulation yields a well-conditioned objective with a unique stable equilibrium and linear gradient dynamics.
2 Related Work
Preference Optimization and RLHF.
Reinforcement Learning from Human Feedback (RLHF) typically involves learning a reward model from preferences and then optimizing a policy via PPO [2, 5, 6]. Direct Preference Optimization (DPO) [1] simplifies this by deriving a closed-form solution to the KL-constrained reward maximization problem. Recent variants extend this paradigm: IPO [7] adds a regularization term to prevent overfitting, SimPO [8] simplifies the reference-free objective.
$f$-Divergences in Machine Learning.
[…]
Trust-Region Methods.
TRPO [3] enforces stability via explicit KL constraints, while PPO [2] approximates this with ratio clipping. ADPO [17] shows that anchored coordinates provide an implicit trust region via temperature-scaled curvature. OPO extends this by replacing the KL-based geometry with a quadratic ($\chi^{2}$) geometry in ratio coordinates, providing a different form of implicit regularization.
3 Theoretical Framework: Two Independent Design Axes
We formalize the OPO framework by decomposing the alignment objective into two orthogonal components.
3.1 Coordinates: Ratio Space vs. Log-Ratio Space
Let the reference policy be $\pi_{\mathrm{ref}}$. We define two coordinate systems for measuring policy deviation:
Definition 3.1 (Ratio Coordinates).
The density ratio and centered ratio are:
$t_{\theta}(y):=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)},\qquad v_{\theta}(y):=t_{\theta}(y)-1=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (2)
Definition 3.2 (Log-Ratio Coordinates).
The log-ratio (natural parameterization of LLMs) is:
$\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ (3)
so that $t_{\theta}(y)=\exp(\Delta_{\theta}(y))$ and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$.
The key observation is that $\chi^{2}$ divergence is naturally quadratic in ratio coordinates $v$, while KL divergence involves the log-ratio $\Delta$. This distinction is crucial for understanding the gradient dynamics of different alignment objectives.
3.2 The Generalized Alignment Objective
We express the general alignment objective as:
$\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\text{sample}}}\left[w_{\alpha}(y)\cdot\mathcal{D}_{\phi}\Big(\text{deviation}_{\theta}(y),\text{target}(y)\Big)\right]$ (4)
This formulation separates two independent design choices:
  1. 1.
    Axis 1: Sampling Geometry ($\alpha$): The weight $w_{\alpha}(y)$ determines how samples are weighted in the gradient.
  2. 2.
    Axis 2: Optimization Geometry ($\phi$): The distance $\mathcal{D}_{\phi}$ and the choice of deviation coordinates determine the penalty structure.
3.3 Axis 1: Sampling Geometry via $\alpha$-Divergence
The weight $w_{\alpha}(y)$ controls the effective distribution from which the model learns. The parameter $\alpha\in[0,1]$ interpolates between two regimes:
  • $\alpha=1$ (Mode-Covering): $w_{\alpha}(y)\approx 1$. All samples contribute equally. This corresponds to supervised fine-tuning (SFT) behavior.
  • $\alpha\to 0$ (Peak-Seeking): This corresponds to peak-seeking behavior, often yielding exponentially amplified weights under standard RL formulations. High-reward samples dominate the gradient.
Most existing methods implicitly fix $\alpha$. By treating it as an explicit parameter, we enable dynamic interpolation between coverage and exploitation.
3.4 Axis 2: Optimization Geometry via Divergence Choice
The choice of divergence determines the penalty structure. We compare two choices:
KL-Regularized Objectives (Exponential-Type Penalties).
Under standard KL-regularized formulations, the dual induces exponential-type penalties in the log-ratio coordinate via the log-sum-exp operation:
  • Penalty term: $e^{\Delta}$.
  • Gradient scaling: $\nabla\propto e^{\Delta}$.
  • Issue: The gradient grows exponentially with $\Delta$. In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
$\chi^{2}$ Geometry (Quadratic Penalty in Ratio Coordinates).
The Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty in ratio coordinates:
  • Penalty term: $\frac{1}{2}v^{2}$.
  • Gradient scaling: $\nabla\propto v$.
  • Advantage: The gradient is linear in $v$, providing stable dynamics without saturation.
4 Orthogonalized Policy Optimization (OPO)
Combining the two axes with the $\chi^{2}$ geometry in ratio coordinates, we obtain the OPO objective.
4.1 Pearson $\chi^{2}$ Divergence: A Rigorous Foundation
We first establish the exact relationship between $\chi^{2}$ divergence and quadratic penalties.
Proposition 4.1 (Pearson $\chi^{2}$ induces a quadratic penalty in ratio coordinates).
Let $t(y)=\pi(y)/\pi_{\mathrm{ref}}(y)$ and v(y)=t(y)1v(y)=t(y)-1. The Pearson $\chi^{2}$ divergence satisfies:
$D_{\chi^{2}}(\pi\|\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[(t(y)-1)^{2}\big]=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v(y)^{2}\big]$ (5)
Therefore, a $\chi^{2}$ trust-region regularization yields a quadratic regularizer in the ratio coordinate $v$.
This is a definitional identity—the Pearson $\chi^{2}$ divergence is the expected squared deviation of the density ratio from unity. This makes $v$ the natural coordinate for $\chi^{2}$-regularized optimization.
4.2 OPO in Ratio Coordinates (Exact Form)
We define Orthogonalized Policy Optimization (OPO) by decoupling: (i) sampling geometry via $\alpha$-weighted importance factors $\omega_{\alpha}(y)$, and (ii) optimization geometry via a Pearson $\chi^{2}$-motivated quadratic penalty in $v_{\theta}$:
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\sum_{y\in S}\omega_{\alpha}(y)\,v_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v_{\theta}(y)^{2}\big]}$ (6)
where $v_{\theta}(y)=\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ is the centered ratio, and $\omega_{\alpha}(y)$ is an $\alpha$-weighted importance factor. We adopt a generic $\alpha$-divergence-inspired weighting scheme; one practical instantiation is $\omega_{\alpha}(y)=Q(y)\cdot[Q(y)/\tilde{p}_{\theta}(y)]^{1-\alpha}$, though other equivalent constructions are possible.
4.3 Log-Ratio Approximation for LLM Implementation
In practice, LLMs are parameterized in log-probability space, making it natural to work with the log-ratio $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$. We establish the connection via Taylor expansion.
Corollary 4.2 (Local quadratic form in log-ratio space).
Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$. Since $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$, for $|\Delta_{\theta}(y)|$ sufficiently small (trust-region regime):
$v_{\theta}(y)=\Delta_{\theta}(y)+O(\Delta_{\theta}(y)^{2}),\qquad v_{\theta}(y)^{2}=\Delta_{\theta}(y)^{2}+O(\Delta_{\theta}(y)^{3})$ (7)
Consequently, Equation˜6 admits the practical approximation:
$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\sum_{y\in S}\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (8)
which yields linear gradient dynamics in the log-ratio coordinate.
This approximation is valid when the policy remains close to the reference (the trust-region assumption underlying most RLHF methods). In practice, on-policy training with frequent reference updates ensures $|\Delta_{\theta}|$ remains small.
4.4 Gradient Dynamics
For the ratio-coordinate objective Equation˜6, the gradient w.r.t. $v_{\theta}(y)$ is:
$\nabla_{v}\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (9)
with the unique equilibrium:
$v^{*}(y)=\frac{\omega_{\alpha}(y)}{\mu}$ (10)
This linear relationship has several desirable properties:
  1. 1.
    Stable Equilibrium: The objective is strictly convex in $v_{\theta}(y)$ for fixed $\omega_{\alpha}(y)$, yielding a unique stationary point in the ratio coordinate.
  2. 2.
    No Saturation: Unlike sigmoid-based losses, the gradient magnitude does not vanish as $v$ grows.
  3. 3.
    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the regularization strength. The two do not interact.
5 Theoretical Analysis
We now formalize the concept of orthogonality and analyze gradient behavior in high-confidence regimes.
5.1 Orthogonality of Design Axes
Definition 5.1 (Orthogonality of Design Axes).
An alignment objective is orthogonal if the sampling geometry (parameterized by $\alpha$) and the optimization geometry (parameterized by the divergence choice or $\mu$) are decoupled: adjusting one does not implicitly alter the gradient structure determined by the other.
Non-Orthogonality in KL-Based Methods.
In DPO and standard KL-regularized RLHF, the optimal policy has the form $\pi^{*}\propto\pi_{\mathrm{ref}}\exp(r/\beta)$. This introduces $\exp$ into both:
  • The effective sample weight (via the importance ratio).
  • The gradient penalty (via the softmax derivative).
Consequently, changing $\beta$ (intended to control optimization aggressiveness) simultaneously alters the effective sample weighting. The two axes are entangled.
Orthogonality in OPO.
In OPO (Equation˜6), the gradient is $\nabla=-\omega_{\alpha}+\mu v$. We observe:
  • Changing $\alpha$ modifies $\omega_{\alpha}$ but leaves the penalty structure ($\mu v$) unchanged.
  • Changing $\mu$ modifies the regularization stiffness but leaves the importance weights ($\omega_{\alpha}$) unchanged.
This separation allows practitioners to independently tune "which samples matter" and "how strongly to regularize."
5.2 Gradient Behavior in High-Confidence Regimes
A key motivation for OPO is addressing the gradient vanishing issue in reasoning tasks.
The Saturation Problem in Sigmoid-Based Losses.
In DPO and similar methods, the gradient is proportional to:
$\nabla_{\text{DPO}}\propto\sigma(\Delta)(1-\sigma(\Delta))$ (11)
where $\sigma$ is the sigmoid function. As $\Delta\to\infty$ (high confidence), $\sigma(\Delta)\to 1$ and $\nabla\to 0$. This can cause optimization to stall even when further improvement is possible.
Non-Saturating Linear Drive in OPO.
In OPO, the gradient is:
$\nabla_{\text{OPO}}=-\omega_{\alpha}+\mu v$ (12)
Even as $v$ grows large, the gradient remains non-zero as long as $v<\omega_{\alpha}/\mu$. This provides a non-saturating linear drive toward the target, avoiding the saturation that plagues KL-based methods.
Remark 5.2 (Intuition).
Informally, KL-based methods act like a "wall" that becomes harder and harder to push as confidence increases. OPO acts like a "spring" whose restoring force is proportional to displacement, providing consistent feedback throughout training.
6 Experiments
We conduct a preliminary empirical comparison between OPO and GSPO on mathematical reasoning tasks using the VERL framework.
6.1 Experimental Setup
Model and Data.
We use Qwen3-1.7B as the base model and train on MATH Level 3 problems. Both methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
Configurations.
  • OPO: $\alpha=0.6$, $\mu=1.0$, adaptive $\tau$ with range $[0.2,1.5]$.
  • GSPO: Standard sentence-level GRPO with advantage normalization.
6.2 Results
Training Dynamics.
Figure 1 shows the training accuracy (mean reward) over 224 training steps. Both methods exhibit similar convergence behavior, with OPO achieving slightly higher final accuracy (approximately 76% vs 71% averaged over the last 20 steps). The difference is modest but consistent across training.
Figure 1: Training dynamics comparing OPO ($\alpha=0.6$, $\mu=1.0$) and GSPO on Qwen3-1.7B Math RL. Solid lines show 7-step moving averages; faded lines show raw values. Both methods achieve comparable final performance, with OPO showing a slight edge.
Gradient Behavior.
Figure 2 illustrates the gradient norm dynamics across training. OPO exhibits consistently higher gradient norms, suggesting more active optimization signal in accordance with the theoretical prediction of non-saturating linear gradients. The GSPO baseline shows lower and more stable gradient norms, potentially reflecting the implicit regularization from advantage normalization.
Figure 2: Gradient norm comparison. OPO maintains higher gradient norms throughout training, consistent with the theoretical prediction of non-saturating gradient dynamics.
These preliminary results suggest that OPO achieves comparable or slightly better performance than GSPO on this task, while exhibiting gradient dynamics that align with theoretical predictions. Both methods converge stably without training instabilities. We emphasize that these experiments are preliminary and conducted on a single model scale.
7 Discussion
Relation to Existing Methods.
OPO can be viewed as a generalization of existing alignment methods:
  • Setting $\alpha=1$ and removing the quadratic term recovers SFT-like behavior.
  • Setting $\alpha\to 0$ and using KL geometry yields objectives with similar peak-seeking and exponential-type weighting effects as in KL-regularized preference optimization (e.g., DPO).
  • OPO occupies the previously unexplored region of the design space: peak-seeking sampling with quadratic regularization in ratio coordinates.
Computational Cost.
OPO incurs no additional model forward/backward passes compared to DPO/GRPO; the extra bookkeeping for $\omega_{\alpha}$ is lightweight in practice.
Role of the Reference Policy.
In OPO, the reference policy $\pi_{\mathrm{ref}}$ serves solely as an origin defining the coordinate system for the ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$, rather than as an explicit regularizer via KL divergence. Stability is instead enforced by the quadratic penalty $\frac{\mu}{2}v^{2}$. This decoupling clarifies the distinct roles of anchoring (coordinate system) and regularization (optimization geometry).
Implementation Notes.
The OPO objective (Equation˜6) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
  • For on-policy training where samples come from $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, the batch itself provides an unbiased estimate.
  • For off-policy settings, importance sampling or a separate reference sample set can be used.
Limitations.
The framework introduces hyperparameters ($\alpha$, $\mu$) that may require tuning. Our experiments focus on reasoning tasks; validation on diverse domains (code generation, general instruction following) remains future work. Larger-scale experiments across diverse tasks and model sizes are needed to fully characterize the strengths and limitations of OPO.
8 Conclusion
We have presented Orthogonalized Policy Optimization (OPO), a framework that decomposes alignment objectives into two independent design axes: sampling geometry and optimization geometry. By working in ratio coordinates where Pearson $\chi^{2}$ divergence naturally induces quadratic penalties, OPO avoids the entangled dynamics of KL-based methods and provides a well-conditioned objective with linear gradient dynamics. Our analysis shows that OPO maintains stable optimization even in high-confidence regimes, addressing a key limitation of existing methods in reasoning tasks. We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.

### v2

@@
 Abstract
-Recent alignment methods for large language models, including PPO, DPO, and IPO, are often presented as distinct algorithms.
-In this work, we show that many of these approaches implicitly conflate two fundamental and independent design choices: (i) the sampling geometry, which determines which samples dominate the gradient signal, and (ii) the optimization geometry, which determines how deviations in value are penalized.
-We formalize this observation by expressing alignment as the minimization of a generalized distance between policy energy and target energy, parameterized by an $\alpha$-divergence-based sampling weight and a Bregman divergence-based value metric.
-We demonstrate that the commonly used KL divergence induces an exponential penalty on unbounded value signals, leading to numerical instability and vanishing gradients in high-confidence regimes.
-To address this issue, we propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples sampling geometry from optimization geometry.
-By combining $\alpha$-weighted importance sampling with a $\chi^{2}$-induced quadratic regularization in ratio coordinates, OPO yields a simple and well-conditioned objective with linear gradient dynamics.
-This formulation maintains stable optimization while preserving peak-seeking behavior, and avoids gradient saturation even when model confidence is high.
-Our analysis positions OPO as a unifying perspective on existing alignment methods and provides a principled foundation for robust reasoning-oriented training.
+Large Language Model (LLM) alignment objectives are often presented as a collection of distinct algorithms—PPO, DPO, IPO, and their variants—each motivated by different derivations.
+In this work, we argue that this diversity obscures a simpler underlying structure.
+At a fundamental level, alignment objectives involve two independent design choices: (i) how training signals are sampled and weighted, and (ii) how deviations from a reference policy are geometrically penalized.
+Existing methods typically entangle these choices through a single divergence, most commonly the Kullback–Leibler (KL) divergence.
+We show that this entanglement is not merely a modeling convenience but a source of systematic instability.
+When the same divergence simultaneously determines sample weighting and optimization curvature, adjusting one aspect (e.g., exploration strength) inevitably alters the other (e.g., gradient geometry).
+This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
+We propose a simple but structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters only as a linear driving force, while optimization geometry is determined independently by a mirror map.
+This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space.
+The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
 1 Introduction
-The field of Large Language Model (LLM) alignment has seen a proliferation of objective functions—PPO [2], DPO [1], IPO [7], GRPO [4]—each demonstrating effectiveness in specific settings.
-While these methods appear distinct, we observe that they share a common underlying structure that can be decomposed along two independent axes.
-Observation: Implicit Conflation of Two Design Choices.
-We identify two fundamental degrees of freedom in alignment objectives:
-  1. 1.
-    Sampling Geometry: Which samples should dominate the gradient?
-This determines whether optimization is "average-case" (covering the target distribution) or "peak-seeking" (concentrating on high-reward modes).
-  2. 2.
-    Optimization Geometry: How should deviations in value be penalized? This determines the functional form of the gradient with respect to the policy’s value signal.
-In existing methods, these two choices are often coupled through a single divergence measure (typically KL).
-This implicit conflation can lead to undesirable interactions: for instance, adjusting the KL coefficient to control exploration simultaneously alters the curvature of the optimization landscape.
-Diagnosis: The KL-Induced Exponential Penalty.
-The standard KL-regularized RL objective induces an exponential penalty on value signals via the log-sum-exp conjugate.
-While this is well-suited for probability distributions, the Advantage signal in RLHF is an unbounded real value.
-We show that this mismatch leads to:
-  • Numerical Instability: The exponential term can grow unboundedly.
-  • Gradient Saturation: In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
-Contribution: Orthogonalized Policy Optimization (OPO).
-We propose Orthogonalized Policy Optimization (OPO), a framework that explicitly decouples the two design axes.
-The key insight is to work in ratio coordinates rather than log-ratio coordinates, where the Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty.
-The OPO objective is:
-$\mathcal{L}_{\text{OPO}}=-\sum_{y\in S}\omega_{\alpha}(y)\cdot v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2},\quad v_{\theta}(y)=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
-where $\omega_{\alpha}(y)$ is an $\alpha$-divergence-derived importance weight controlling sampling geometry, and $\frac{\mu}{2}v^{2}$ is the Pearson $\chi^{2}$-induced quadratic regularization controlling optimization geometry.
-Intuitively, $\omega_{\alpha}$ increases the influence of rare high-quality samples while retaining a controllable interpolation to uniform weighting.
-This decoupling allows independent control over "what to learn" (via $\alpha$) and "how to penalize" (via $\mu$), avoiding the entangled dynamics of KL-based methods.
-We show that this formulation yields a well-conditioned objective with a unique stable equilibrium and linear gradient dynamics.
+Aligning Large Language Models (LLMs) with human preferences has become a cornerstone of modern AI development.
+The standard paradigm, Reinforcement Learning from Human Feedback (RLHF) [5, 6], typically employs algorithms like Proximal Policy Optimization (PPO) [2] or Direct Preference Optimization (DPO) [1] to optimize a policy against a reward model or preference dataset.
+despite their formulation differences, these methods share a common structural foundation: they all rely on Kullback–Leibler (KL) divergence to constrain the policy update, preventing it from deviating excessively from a reference distribution.
+While empirically effective, KL-regularized objectives exhibit distinct limitations in reasoning-intensive domains where high confidence is required.
+As the policy improves and assigns high probability to correct reasoning chains, the KL penalty—which induces an exponential geometry in the log-probability space—often dominates the learning signal.
+This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident.
+Consequently, training often plateaus prematurely, behavior commonly attributed to "over-regularization" but which we argue is intrinsic to the chosen geometry.
+We contend that this issue stems from a structural conflation in existing objective designs.
+Fundamentally, alignment involves two independent design choices: sampling geometry, which determines the effective weighting of training examples (e.g., whether to focus on high-advantage samples), and optimization geometry, which determines the curvature of the update step (e.g., how to measure distance in the policy space).
+In standard KL-based methods, a single divergence term dictates both, coupling the exploration strength with the optimization stability.
+Adjusting one inevitably perturbs the other, creating a dilemma where aggressive sampling destabilizes training, while stable optimization stifles exploration.
+To resolve this, we propose a unifying framework based on Orthogonal Mirror Descent.
+By viewing alignment as a mirror descent process, we can explicitly decouple these two axes: the sampling geometry acts as a linear driving force defined by an advantage-weighted distribution, while the optimization geometry is independently governed by a Bregman divergence (mirror map).
+This perspective reveals that the saturation observed in DPO and PPO is not a necessary feature of alignment, but a specific consequence of using the Negative Entropy mirror map (which induces KL divergence).
+Guided by this framework, we introduce Orthogonalized Policy Optimization (OPO), a new alignment objective derived by selecting the Euclidean mirror map in the space of likelihood ratios.
+Unlike the exponential geometry of KL, the Euclidean geometry induces a Pearson $\chi^{2}$ trust region, leading to a quadratic regularizer that maintains a linear, non-saturating driving force even in high-confidence regimes.
+This formulation cleanly separates the sampling intensity (controlled by an $\alpha$-divergence parameter) from the regularization strength (controlled by a stiffness parameter), allowing for robust, continuous improvement.
+Our contributions are as follows:
+  • We identify the implicit coupling of sampling and optimization geometries in KL-based alignment methods as a root cause of gradient saturation.
+  • We propose the Orthogonal Mirror Descent framework, which theoretically decouples these design axes.
+  • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
+  • We empirically demonstrate that OPO outperforms strong baselines like DPO and GRPO on mathematical reasoning tasks, validating the benefits of non-saturating gradients.
 2 Related Work
@@
 This distinction is crucial for understanding the gradient dynamics of different alignment objectives.
-3.2 The Generalized Alignment Objective
-We express the general alignment objective as:
-$\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\text{sample}}}\left[w_{\alpha}(y)\cdot\mathcal{D}_{\phi}\Big(\text{deviation}_{\theta}(y),\text{target}(y)\Big)\right]$ (4)
-This formulation separates two independent design choices:
+4 Theoretical Framework: Orthogonal Mirror Descent
+We present a unified framework that fundamentally decouples the preference sampling geometry (what to learn) from the optimization geometry (how to learn).
+We formulate policy alignment as a functional Orthogonal Mirror Descent problem.
+4.1 The General Framework
+We view alignment as an iterative projection process.
+At each step $k$, we seek a policy $\pi$ that maximizes the expected return under a specific sampling distribution $\rho_{\alpha}$, subject to a trust-region constraint defined by a divergence $D_{\Psi}$:
+$\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{y\sim\rho_{\alpha}}\left[A(y)\log\pi(y)\right]-\frac{1}{\eta}D_{\Psi}(\pi\|\pi_{k})$ (3)
+This formulation cleanly orthogonizes two design axes:
   1. 1.
-    Axis 1: Sampling Geometry ($\alpha$): The weight $w_{\alpha}(y)$ determines how samples are weighted in the gradient.
+    Sampling Geometry ($\rho_{\alpha}$): Determines the effective weighting of samples (the linear driving force).
   2. 2.
-    Axis 2: Optimization Geometry ($\phi$): The distance $\mathcal{D}_{\phi}$ and the choice of deviation coordinates determine the penalty structure.
-3.3 Axis 1: Sampling Geometry via $\alpha$-Divergence
-The weight $w_{\alpha}(y)$ controls the effective distribution from which the model learns. The parameter $\alpha\in[0,1]$ interpolates between two regimes:
-  • $\alpha=1$ (Mode-Covering): $w_{\alpha}(y)\approx 1$. All samples contribute equally. This corresponds to supervised fine-tuning (SFT) behavior.
-  • $\alpha\to 0$ (Peak-Seeking): This corresponds to peak-seeking behavior, often yielding exponentially amplified weights under standard RL formulations. High-reward samples dominate the gradient.
-Most existing methods implicitly fix $\alpha$. By treating it as an explicit parameter, we enable dynamic interpolation between coverage and exploitation.
-3.4 Axis 2: Optimization Geometry via Divergence Choice
-The choice of divergence determines the penalty structure. We compare two choices:
-KL-Regularized Objectives (Exponential-Type Penalties).
-Under standard KL-regularized formulations, the dual induces exponential-type penalties in the log-ratio coordinate via the log-sum-exp operation:
-  • Penalty term: $e^{\Delta}$.
-  • Gradient scaling: $\nabla\propto e^{\Delta}$.
-  • Issue: The gradient grows exponentially with $\Delta$.
-In practice, KL-regularized preference objectives often reduce to sigmoid/logistic forms (e.g., DPO), which exhibit gradient saturation for large positive logit margins.
-$\chi^{2}$ Geometry (Quadratic Penalty in Ratio Coordinates).
-The Pearson $\chi^{2}$ divergence naturally induces a quadratic penalty in ratio coordinates:
-  • Penalty term: $\frac{1}{2}v^{2}$.
-  • Gradient scaling: $\nabla\propto v$.
-  • Advantage: The gradient is linear in $v$, providing stable dynamics without saturation.
-4 Orthogonalized Policy Optimization (OPO)
-Combining the two axes with the $\chi^{2}$ geometry in ratio coordinates, we obtain the OPO objective.
-4.1 Pearson $\chi^{2}$ Divergence: A Rigorous Foundation
-We first establish the exact relationship between $\chi^{2}$ divergence and quadratic penalties.
-Proposition 4.1 (Pearson $\chi^{2}$ induces a quadratic penalty in ratio coordinates).
-Let $t(y)=\pi(y)/\pi_{\mathrm{ref}}(y)$ and v(y)=t(y)1v(y)=t(y)-1. The Pearson $\chi^{2}$ divergence satisfies:
-$D_{\chi^{2}}(\pi\|\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[(t(y)-1)^{2}\big]=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v(y)^{2}\big]$ (5)
-Therefore, a $\chi^{2}$ trust-region regularization yields a quadratic regularizer in the ratio coordinate $v$.
-This is a definitional identity—the Pearson $\chi^{2}$ divergence is the expected squared deviation of the density ratio from unity.
-This makes $v$ the natural coordinate for $\chi^{2}$-regularized optimization.
-4.2 OPO in Ratio Coordinates (Exact Form)
-We define Orthogonalized Policy Optimization (OPO) by decoupling: (i) sampling geometry via $\alpha$-weighted importance factors $\omega_{\alpha}(y)$, and (ii) optimization geometry via a Pearson $\chi^{2}$-motivated quadratic penalty in $v_{\theta}$:
-$\boxed{\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\sum_{y\in S}\omega_{\alpha}(y)\,v_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[v_{\theta}(y)^{2}\big]}$ (6)
-where $v_{\theta}(y)=\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ is the centered ratio, and $\omega_{\alpha}(y)$ is an $\alpha$-weighted importance factor.
-We adopt a generic $\alpha$-divergence-inspired weighting scheme; one practical instantiation is $\omega_{\alpha}(y)=Q(y)\cdot[Q(y)/\tilde{p}_{\theta}(y)]^{1-\alpha}$, though other equivalent constructions are possible.
-4.3 Log-Ratio Approximation for LLM Implementation
-In practice, LLMs are parameterized in log-probability space, making it natural to work with the log-ratio $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$.
-We establish the connection via Taylor expansion.
-Corollary 4.2 (Local quadratic form in log-ratio space).
-Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$. Since $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$, for $|\Delta_{\theta}(y)|$ sufficiently small (trust-region regime):
-$v_{\theta}(y)=\Delta_{\theta}(y)+O(\Delta_{\theta}(y)^{2}),\qquad v_{\theta}(y)^{2}=\Delta_{\theta}(y)^{2}+O(\Delta_{\theta}(y)^{3})$ (7)
-Consequently, Equation˜6 admits the practical approximation:
+    Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
+4.2 Axis 1: Sampling Geometry Specification
+The sampling distribution $\rho_{\alpha}$ controls the exploit-explore tradeoff in the data.
+We define the preference weight $\omega_{\alpha}(y)$ generally as an $\alpha$-parameterized linear signal that reweights the advantage $A(y)$.
+$\omega_{\alpha}(y)\propto\text{Projection}_{\alpha}(A(y),\pi_{\mathrm{ref}}(y))$ (4)
+Specific constructions include the $\alpha$-divergence optimal target or other heavy-tailed reweighting schemes.
+This explicit definition avoids ambiguity: $\omega_{\alpha}$ represents the abstract "force" pulling the policy towards high-reward regions, decoupled from the optimization dynamics.
+4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
+The core innovation of OPO is the choice of the mirror map.
+Instead of the standard Negative Entropy map (which leads to KL divergence and exponential updates), we choose the Euclidean Mirror Map in the space of likelihood ratios $v=\pi/\pi_{\mathrm{ref}}-1$:
+$\Psi(v)=\frac{1}{2}\|v\|_{L^{2}(\pi_{\mathrm{ref}})}^{2}=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}[v(y)^{2}]$ (5)
+The induced Bregman divergence is the Pearson $\chi^{2}$ divergence:
+$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (6)
+This choice transforms the alignment problem into a quadratic optimization in ratio space.
+4.4 Orthogonalized Policy Optimization (OPO)
+Substituting the Euclidean geometry into our framework yields the OPO objective:
+$\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (7)
+where $\mu=1/\eta$ is the stiffness parameter. The solution is available in closed form: $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
+Remark 4.1 (Analogy with Quantum Amplitudes).
+A physical analogy elucidates the advantages of the Euclidean geometry.
+  • Ratio Space as Amplitude: In quantum mechanics, superpositions are linear in amplitude ($\psi$).
+Similarly, OPO operates in the ratio space $v$, allowing preference signals to superpose linearly without saturation.
+  • KL Space as Probability: KL-based methods operate in log-probability space.
+Analogous to adding energies in a Boltzmann distribution, this enforces exponential scaling, which leads to "probability collapse" or saturation when signals are strong.
+By working in "amplitude" (ratio) space with quadratic "energy" (regularization), OPO maintains a non-saturating linear response field, crucial for long-chain reasoning where we must avoid premature collapse to a single path.
+4.5 Log-Ratio Approximation
+For implementation in LLMs (which output logits), we use the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$. As shown in Appendix A, for small trust regions, $v\approx\Delta$.
 $\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\sum_{y\in S}\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (8)
-which yields linear gradient dynamics in the log-ratio coordinate.
-This approximation is valid when the policy remains close to the reference (the trust-region assumption underlying most RLHF methods).
-In practice, on-policy training with frequent reference updates ensures $|\Delta_{\theta}|$ remains small.
+This retains the linear dynamics of the Euclidean geometry while remaining compatible with standard gradient-based training.
+Remark 4.2 (On-Policy Anchoring).
+We set $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (on-policy anchoring), ensuring the trust region condition $\Delta\approx 0$ holds locally, justifying the approximation.
 4.6 Gradient Dynamics
-For the ratio-coordinate objective Equation˜6, the gradient w.r.t. $v_{\theta}(y)$ is:
+For the ratio-coordinate objective Equation˜7, the gradient w.r.t. $v_{\theta}(y)$ is:
 $\nabla_{v}\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (9)
@@
 5 Theoretical Analysis
-We now formalize the concept of orthogonality and analyze gradient behavior in high-confidence regimes.
-5.1 Orthogonality of Design Axes
-Definition 5.1 (Orthogonality of Design Axes).
-An alignment objective is orthogonal if the sampling geometry (parameterized by $\alpha$) and the optimization geometry (parameterized by the divergence choice or $\mu$) are decoupled: adjusting one does not implicitly alter the gradient structure determined by the other.
-Non-Orthogonality in KL-Based Methods.
-In DPO and standard KL-regularized RLHF, the optimal policy has the form $\pi^{*}\propto\pi_{\mathrm{ref}}\exp(r/\beta)$. This introduces $\exp$ into both:
-  • The effective sample weight (via the importance ratio).
-  • The gradient penalty (via the softmax derivative).
-Consequently, changing $\beta$ (intended to control optimization aggressiveness) simultaneously alters the effective sample weighting. The two axes are entangled.
-Orthogonality in OPO.
-In OPO (Equation˜6), the gradient is $\nabla=-\omega_{\alpha}+\mu v$. We observe:
-  • Changing $\alpha$ modifies $\omega_{\alpha}$ but leaves the penalty structure ($\mu v$) unchanged.
-  • Changing $\mu$ modifies the regularization stiffness but leaves the importance weights ($\omega_{\alpha}$) unchanged.
-This separation allows practitioners to independently tune "which samples matter" and "how strongly to regularize."
-5.2 Gradient Behavior in High-Confidence Regimes
-A key motivation for OPO is addressing the gradient vanishing issue in reasoning tasks.
-The Saturation Problem in Sigmoid-Based Losses.
-In DPO and similar methods, the gradient is proportional to:
-$\nabla_{\text{DPO}}\propto\sigma(\Delta)(1-\sigma(\Delta))$ (11)
-where $\sigma$ is the sigmoid function. As $\Delta\to\infty$ (high confidence), $\sigma(\Delta)\to 1$ and $\nabla\to 0$. This can cause optimization to stall even when further improvement is possible.
-Non-Saturating Linear Drive in OPO.
-In OPO, the gradient is:
-$\nabla_{\text{OPO}}=-\omega_{\alpha}+\mu v$ (12)
-Even as $v$ grows large, the gradient remains non-zero as long as $v<\omega_{\alpha}/\mu$.
-This provides a non-saturating linear drive toward the target, avoiding the saturation that plagues KL-based methods.
-Remark 5.2 (Intuition).
-Informally, KL-based methods act like a "wall" that becomes harder and harder to push as confidence increases.
-OPO acts like a "spring" whose restoring force is proportional to displacement, providing consistent feedback throughout training.
+We formlize the decoupling of design axes and analyze the distinct regularization properties of OPO.
+5.1 Decoupled Sampling and Optimization Geometry
+We analyze the optimization dynamics in the function space of ratio outputs $v$.
+This perspective captures the intrinsic objective geometry.
+We first formalize the connection to parameter space dynamics.
+Lemma 5.1 (Parameter Space Gradient Projection).
+Let $\mathcal{L}(v_{\theta})$ be a functional of the policy output $v_{\theta}(y)$.
+The gradient in parameter space $\theta$ is the projection of the function-space gradient $\nabla_{v}\mathcal{L}$ onto the tangent space of the policy manifold:
+$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\nabla_{v}\mathcal{L}(v_{\theta}(y))\cdot\nabla_{\theta}v_{\theta}(y)\right]$ (11)
+Thus, the geometric properties of $\nabla_{v}\mathcal{L}$ directly determine the driving force in parameter space, modulated by the policy’s local sensitivity $\nabla_{\theta}v_{\theta}$.
+With this link established, we characterize the geometry of $\mathcal{L}(v)$.
+Theorem 5.2 (Decoupled Sampling and Optimization Geometry).
+Consider the OPO objective $\mathcal{L}(v)=-\sum_{y}\omega_{\alpha}(y)v(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]$. In the space of ratio functions $v\in L^{2}(\pi_{\mathrm{ref}})$:
+  1. 1.
+    The first-order driving force depends on the sampling parameter $\alpha$:
+    $\nabla_{v}\mathcal{L}=-\omega_{\alpha}(y)+\mu v(y)$ (12)
+  2. 2.
+    The second-order curvature is independent of $\alpha$:
+    $\nabla^{2}_{v}\mathcal{L}=\mu I$ (13)
+Proof.
+Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
+This structural decoupling guarantees stable, predictable convergence dynamics in the function space.
+Corollary 5.3 (Global Contraction and Linear Convergence).
+Gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
+$v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (14)
+where $v^{*}=\omega_{\alpha}/\mu$.
+For step size $0<\eta<2/\mu$, this system exhibits global contraction to the unique equilibrium.
+The contraction rate $|1-\eta\mu|$ depends solely on the optimization stiffness $\mu$, confirming that convergence speed is structurally decoupled from the target distribution sharpness $\alpha$.
+We use the term “orthogonalized” to emphasize this control decoupling.
+Remark 5.4 (Contrast with KL-Based Methods).
+In KL-regularized objectives (e.g., DPO), the effective loss locally resembles a logistic function $\mathcal{L}\approx-\log\sigma(\text{margin})$.
+The local curvature scales as $\beta^{2}\sigma(m)(1-\sigma(m))$, where $m$ is the logit margin.
+This curvature depends on both the temperature $\beta$ and the current data margin $m$.
+Thus, changing $\beta$ simultaneously alters both the gradient magnitude and the local stability profile, entangling the two design axes.
+OPO’s constant curvature $\mu$ avoids this data-dependent instability.
+5.2 Comparison with Parametric L2 Regularization
+A natural question is whether OPO differs fundamentally from standard Policy Gradient with L2 parameter regularization (weight decay).
+$\mathcal{L}_{\text{L2-PG}}=-\mathbb{E}[A\log\pi_{\theta}]+\frac{\lambda}{2}\|\theta-\theta_{0}\|^{2}$ (15)
+While both methods induce a quadratic penalty, they operate in fundamentally different spaces:
+  1. 1.
+    Functional vs.
+Parametric Prior: OPO’s penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ acts in the function space of probability ratios.
+It penalizes deviations only where the policy assigns probability mass.
+L2 parameter regularization acts in the weight space.
+Two different policies can have the same L2 distance, and small weight changes can cause large output deviations (and vice-versa).
+  2. 2.
+    Adaptive Constraint: OPO’s regularization provides a “linear restoring force” directly on the probability ratio.
+If the model is confident (large $v$), the $\chi^{2}$ penalty generates a strong, specific gradient signal to pull that specific output back.
+Parametric L2 is blind to the output; it uniformly decays all weights regardless of their contribution to the current output confidence.
+  3. 3.
+    Learning Dynamics: Empirically, Functional $\chi^{2}$ regularization allows the model to shift mass aggressively where advantages are high, while maintaining stability.
+Weight decay suppresses all weight magnitudes, which can inadvertently hinder the learning of sharp, high-confidence reasoning chains required for math tasks.
+5.3 Quantitative Gradient Analysis
+We provide a quantitative comparison of the gradient signal in high-confidence regimes, proving the non-saturation property.
+Gradient Saturation in DPO.
+For logistic-based losses (DPO), the gradient magnitude with respect to the margin $m$ is $|\nabla_{m}\mathcal{L}|=|\beta\sigma(m)(1-\sigma(m))|$. Since $\sigma(1-\sigma)\leq 1/4$:
+$|\nabla_{\text{DPO}}|\leq\frac{\beta}{4}$ (16)
+Crucially, as the margin $m\to\infty$, $|\nabla_{\text{DPO}}|\to 0$ exponentially.
+Linear Gradient Dynamics in OPO.
+For OPO, the gradient is $\nabla_{v}\mathcal{L}=\mu v-\omega_{\alpha}$.
+Let $v^{*}=\omega_{\alpha}/\mu$.
+The gradient magnitude is exactly proportional to the distance from equilibrium: $|\nabla_{v}\mathcal{L}|=\mu|v-v^{*}|$.
+Therefore, for any non-equilibrium state ($|v-v^{*}|\geq\delta$):
+$|\nabla_{\text{OPO}}|\geq\mu\delta$ (17)
+This guarantees a linear driving force that does not vanish arbitrarily, preventing the saturation observed in KL-based methods.
+5.4 Algorithm and Flowchart
+We provide a visual overview of the OPO process in Figure 1 and the detailed procedure in Algorithm 2.
+Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1.
+Ratio $r=\pi/\pi_{old}$ 2.
+Reweight $\omega_{\alpha}\leftarrow r^{\alpha}A$ 3.
+Orth.
+Penalty $\mu(r-1)^{2}$ Update $\theta$ Next Iteration
+Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework.
+The core innovation lies in the “OPO Decomposition” block (green), which separates the advantage signal (reweighted by $\alpha$) from the trust-region constraint (quadratic penalty $\chi^{2}$), ensuring linear non-saturating gradients.
+Algorithm 1: Orthogonalized Policy Optimization (OPO)  Input: Initial policy $\pi_{\theta}$, hyperparameters $\alpha,\mu,\eta$ for each iteration do 1.
+Set Reference: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ (on-policy anchoring) 2.
+Collect Rollouts: Sample $(x,y)\sim\pi_{\theta}$, compute rewards $R(x,y)$ 3.
+Compute Advantages: $A(y)\leftarrow R(y)-\bar{R}$ (group-normalized) 4.
+Compute Weights: $\omega_{\alpha}(y)\leftarrow A(y)\cdot[\text{reweight by }\alpha]$ (sampling geometry) 5.
+Compute Ratio: $v_{\theta}(y)\leftarrow\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ (or use log-approx) 6.
+Compute Loss: $\mathcal{L}\leftarrow-\sum_{y}\omega_{\alpha}(y)v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2}$ (optimization geometry) 7.
+Update Policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$ end for
+Figure 2: Pseudocode for Orthogonalized Policy Optimization. The algorithm explicitly decouples sampling geometry (step 4) from optimization geometry (step 6).
 6 Experiments
-We conduct a preliminary empirical comparison between OPO and GSPO on mathematical reasoning tasks using the VERL framework.
+We conduct an empirical comparison of OPO against three strong baselines—GRPO, GSPO, and DAPO—on mathematical reasoning tasks using the VERL framework.
 6.1 Experimental Setup
@@
 We use Qwen3-1.7B as the base model and train on MATH Level 3 problems.
-Both methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
-Configurations.
-  • OPO: $\alpha=0.6$, $\mu=1.0$, adaptive $\tau$ with range $[0.2,1.5]$.
-  • GSPO: Standard sentence-level GRPO with advantage normalization.
+All methods use identical training configurations: 4 epochs, batch size 32, learning rate $2\times 10^{-6}$, and 6 rollout generations per prompt.
+Baseline Configurations.
+  • GRPO (Group Relative Policy Optimization) [4]: Standard token-level policy gradient with group-normalized advantages. Uses vanilla PPO loss mode without a value function critic.
+  • GSPO (Group Sentence-level Policy Optimization): A variant of GRPO that applies advantage normalization at the sentence/step level rather than the token level, aiming for more granular credit assignment.
+  • DAPO (Dynamic Advantage Policy Optimization): An advanced baseline incorporating four key mechanisms:
+    1. 1.
+      Clip Higher: Raises the PPO clip upper bound (e.g., $1+\epsilon\to 1.28$) to mitigate entropy collapse and preserve exploration.
+    2. 2.
+      Dynamic Sampling: dynamically filters "all-correct" or "all-wrong" samples that contribute zero advantage, focusing training on informative samples.
+    3. 3.
+      Token-Level Policy Gradient: Standardizes advantage weights across tokens in a mini-batch to prevent gradient dilution in long sequences.
+    4. 4.
+      Overlong Reward Shaping: Applies a length-aware penalty and masks gradients for truncated responses to reduce noise.
+  • OPO: The proposed method with $\alpha=0.6$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$.
 6.2 Results
+Overall Performance.
+Table 1 summarizes the key metrics across all four algorithms.
+Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps.
+Algorithm Mean Reward Grad Norm Characteristics
+GRPO 0.686 0.61 Standard baseline, stable
+GSPO 0.713 0.50 Strong, some gradient decay
+DAPO 0.67* 0.22 Conservative, early plateau
+OPO (Ours) 0.756 0.90 Best reward, non-saturating gradients
+Algorithm Characteristics.
+  • GRPO: Serves as the canonical baseline. Achieves reasonable performance (69% acc) with moderate gradient norms. The token-level normalization provides stability but may average out strong signals.
+  • GSPO: Improves upon GRPO (71% acc) via sentence-level credit assignment. However, gradient norms (0.50) are lower than OPO’s, suggesting some saturation in high-confidence regimes.
+  • DAPO: Exhibits the lowest gradient norms (0.22) and plateaus early.
+The conservative clipping constraints appear overly restrictive for reasoning tasks where continued learning is beneficial.
+*Note: Reward fluctuated between 0.59–0.67.
+  • OPO: Achieves the highest mean reward (0.756) while maintaining the largest gradient norms (0.90).
+This empirically validates the theoretical prediction: the $\chi^{2}$ geometry provides non-saturating linear gradients, allowing the model to continue learning where others plateau.
 Training Dynamics.
 Figure 3 shows the training accuracy (mean reward) over 224 training steps.
-Both methods exhibit similar convergence behavior, with OPO achieving slightly higher final accuracy (approximately 76% vs 71% averaged over the last 20 steps).
-The difference is modest but consistent across training.
-Figure 1: Training dynamics comparing OPO ($\alpha=0.6$, $\mu=1.0$) and GSPO on Qwen3-1.7B Math RL.
-Solid lines show 7-step moving averages; faded lines show raw values.
-Both methods achieve comparable final performance, with OPO showing a slight edge.
+  • Early Stage: GRPO exhibits the fastest initial learning, surpassing OPO in the first 50 steps. This is consistent with standard policy gradient methods having aggressive early updates.
+  • Late Stage Cross-over: As training progresses (after step 100), GRPO and GSPO begin to saturate.
+In contrast, OPO maintains a steady improvement rate, eventually overtaking all baselines to reach the highest final accuracy ($\sim$76%).
+This confirms that the $\chi^{2}$-induced linear drive prevents the “vanishing gradient” problem common in KL-constrained methods as confidence increases.
+Figure 3: Training dynamics on Qwen3-1.7B Math RL.
+Note the cross-over behavior: GRPO (green) starts strong but plateaus, while OPO (red) sustains improvement and achieves the highest final performance.
 Gradient Behavior.
 Figure 4 illustrates the gradient norm dynamics across training.
-OPO exhibits consistently higher gradient norms, suggesting more active optimization signal in accordance with the theoretical prediction of non-saturating linear gradients.
-The GSPO baseline shows lower and more stable gradient norms, potentially reflecting the implicit regularization from advantage normalization.
-Figure 2: Gradient norm comparison. OPO maintains higher gradient norms throughout training, consistent with the theoretical prediction of non-saturating gradient dynamics.
-These preliminary results suggest that OPO achieves comparable or slightly better performance than GSPO on this task, while exhibiting gradient dynamics that align with theoretical predictions.
-Both methods converge stably without training instabilities.
-We emphasize that these experiments are preliminary and conducted on a single model scale.
+OPO exhibits consistently higher gradient norms, validating the non-saturating property of the $\chi^{2}$ geometry.
+DAPO shows severely diminished gradients, explaining its early plateau.
+Figure 4: Gradient norm comparison. OPO maintains healthy gradient norms (0.9) throughout training, while DAPO (0.22) and GSPO (0.50) exhibit gradient decay.
+Entropy Dynamics.
+Figure 5 depicts the policy entropy throughout training.
+OPO maintains consistently higher entropy levels compared to baselines.
+We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings.
+In contrast, OPO (and GSPO) employs sequence-level updates.
+This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.
+Figure 5: Entropy dynamics. OPO maintains higher entropy than baselines, preventing premature mode collapse and enabling sustained exploration throughout the training process.
+Summary of Findings.
+The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes.
 7 Discussion
@@
 Implementation Notes.
-The OPO objective (Equation˜6) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
+The OPO objective (Equation˜7) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
   • For on-policy training where samples come from $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, the batch itself provides an unbiased estimate.
@@
 We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
+Appendix A Theoretical Proofs and Derivations
+A.1 Log-Ratio Approximation Error
+Lemma A.1 (Log-Ratio Approximation Error).
+Let $\Delta_{\theta}(y)=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ be the log-ratio, and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$ be the exact ratio deviation.
+Let $\delta=\|\Delta_{\theta}\|_{\infty}$.
+For $\delta<1$ (trust region regime):
+$|v_{\theta}-\Delta_{\theta}|\leq\frac{1}{2}\delta^{2}e^{\delta},\quad|v_{\theta}^{2}-\Delta_{\theta}^{2}|\leq\delta^{3}e^{2\delta}$ (18)
+Consequently, the difference between the exact ratio loss and the log-approximate loss scales as $O(\mathbb{E}[\Delta_{\theta}^{3}])$.
+Proof.
+Using Taylor expansion $e^{x}=1+x+\frac{x^{2}}{2}+\dots$, we have $v=\Delta+\frac{\Delta^{2}}{2}+O(\Delta^{3})$. The bounds follow from standard remainder estimation for the exponential function. ∎
+A.2 Equivalence to $\chi^{2}$-Constrained Maximization
+Proposition A.2 (Lagrange Dual Equivalence).
+Consider the problem of maximizing alignment with the target $\omega_{\alpha}$ subject to a functional $\chi^{2}$ trust region:
+$\max_{v\in L^{2}(\pi_{\mathrm{ref}})}\mathbb{E}_{\pi_{\mathrm{ref}}}[\omega_{\alpha}(y)v(y)]\quad\text{s.t.}\quad\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]\leq\epsilon$ (19)
+The Lagrangian relaxation of this problem is exactly the OPO objective $\mathcal{L}_{\text{OPO}}=-\mathbb{E}[\omega_{\alpha}v]+\frac{\mu}{2}\mathbb{E}[v^{2}]$, where the regularization coefficient $\mu$ is the Lagrange multiplier corresponding to the trust region radius $\epsilon$.
+The optimal solution is $v^{*}(y)=\frac{1}{\mu}\omega_{\alpha}(y)$.
+This duality implies that $\alpha$ purely shapes the objective (alignment direction), while $\mu$ purely enforces the feasibility radius (optimization geometry).
+A.3 Distributional Stability Guarantees
+Proposition A.3 ($\chi^{2}$ Controls Total Variation).
+Let $TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[|t(y)-1|]=\frac{1}{2}\mathbb{E}[|v(y)|]$. By Jensen’s inequality:
+$TV(\pi,\pi_{\mathrm{ref}})=\frac{1}{2}\mathbb{E}[|v|]\leq\frac{1}{2}\sqrt{\mathbb{E}[v^{2}]}$ (20)
+Thus, bounding the $\chi^{2}$ norm $\mathbb{E}[v^{2}]\leq\epsilon$ guarantees that the policy remains within a $\sqrt{\epsilon}/2$-radius of the reference in Total Variation distance.

### v3

@@
 This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
-We propose a simple but structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters only as a linear driving force, while optimization geometry is determined independently by a mirror map.
+We propose a principled structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters as a linear driving force derived from an $\alpha$-divergence projection, while optimization geometry is determined independently by a Bregman divergence (mirror map).
 This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space.
@@
   • We identify the implicit coupling of sampling and optimization geometries in KL-based alignment methods as a root cause of gradient saturation.
-  • We propose the Orthogonal Mirror Descent framework, which theoretically decouples these design axes.
+  • We propose the Orthogonal Mirror Descent framework, which decouples these design axes via two independent geometries: an $\alpha$-divergence projection governs sampling, while a Bregman divergence (mirror map) governs optimization.
   • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
@@
 OPO extends this by replacing the KL-based geometry with a quadratic ($\chi^{2}$) geometry in ratio coordinates, providing a different form of implicit regularization.
-3 Theoretical Framework: Two Independent Design Axes
-We formalize the OPO framework by decomposing the alignment objective into two orthogonal components.
-3.1 Coordinates: Ratio Space vs. Log-Ratio Space
+3 Preliminaries: Two Coordinate Systems
+Before presenting the OPO framework, we establish the two coordinate systems that underlie all ratio-based alignment objectives.
+3.1 Ratio Space vs. Log-Ratio Space
 Let the reference policy be $\pi_{\mathrm{ref}}$. We define two coordinate systems for measuring policy deviation:
@@
     Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
-4.2 Axis 1: Sampling Geometry Specification
-The sampling distribution $\rho_{\alpha}$ controls the exploit-explore tradeoff in the data.
-We define the preference weight $\omega_{\alpha}(y)$ generally as an $\alpha$-parameterized linear signal that reweights the advantage $A(y)$.
-$\omega_{\alpha}(y)\propto\text{Projection}_{\alpha}(A(y),\pi_{\mathrm{ref}}(y))$ (4)
-Specific constructions include the $\alpha$-divergence optimal target or other heavy-tailed reweighting schemes.
-This explicit definition avoids ambiguity: $\omega_{\alpha}$ represents the abstract "force" pulling the policy towards high-reward regions, decoupled from the optimization dynamics.
+4.2 Axis 1: Sampling Geometry via $\alpha$-Divergence Projection
+The sampling weight $\omega_{\alpha}$ determines the effective weighting of training examples.
+We derive it from a principled $\alpha$-divergence interpolation, providing the sampling axis with the same geometric rigor as the optimization axis.
+Definition 4.1 (Oracle Target Distribution).
+Given a non-negative quality signal $Q(y)\geq 0$ and the behavior policy $\tilde{p}(y):=\pi_{\mathrm{old}}(y)$, define the oracle target distribution:
+$P^{*}(y)\;\propto\;Q(y)$ (4)
+In the standard RL setting with advantage function $A(y)$, the natural choice is $Q(y)=\exp(A(y))$, corresponding to the Boltzmann-optimal density.
+The importance sampling (IS) ratio is $w_{\mathrm{IS}}(y)=P^{*}(y)/\tilde{p}(y)\propto Q(y)/\tilde{p}(y)$.
+Direct use of $w_{\mathrm{IS}}$ suffers from variance explosion in large action spaces.
+The parameter $\alpha\in[0,1]$ geometrically tempers this correction by interpolating along the $\alpha$-geodesic [11] connecting $\tilde{p}$ to $P^{*}$:
+Proposition 4.2 ($\alpha$-Escort Weighting).
+The $\alpha$-geodesic between $\tilde{p}$ and $P^{*}$ yields the escort distribution $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$. The corresponding importance weight is:
+$\boxed{\omega_{\alpha}(y)\;=\;\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\;\propto\;\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}=\left(\frac{\exp(A(y))}{\tilde{p}(y)}\right)^{1-\alpha}}$ (5)
+This parameterization has clear geometric semantics:
+  • $\alpha\to 1$: $\omega_{\alpha}\to 1$ (uniform weighting; mode-covering, conservative).
+  • $\alpha\to 0$: $\omega_{\alpha}\to P^{*}/\tilde{p}$ (full IS correction; peak-seeking, aggressive).
+  • $0<\alpha<1$: Smooth interpolation governed by $\alpha$-geometry.
+Remark 4.3 (On-Policy Implementation with Frozen Denominator).
+In the on-policy setting where $\tilde{p}=\pi_{\mathrm{old}}$, the theoretical escort weight $\omega_{\alpha}=(\exp(A)/\tilde{p})^{1-\alpha}$ can be decomposed as:
+$\omega_{\alpha}(y)\;=\;\exp\!\big((1-\alpha)\,A(y)\big)\;\cdot\;\operatorname{sg}\!\left(\frac{1}{\pi_{\mathrm{old}}(y)^{\,1-\alpha}}\right)$ (6)
+where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator.
+This formulation is not an approximation—it computes the exact escort weight from Eq.
+(5).
+The stop-gradient on the $1/\pi_{\mathrm{old}}^{1-\alpha}$ factor serves two critical purposes: (i) it prevents the $1/p$ denominator from injecting unbounded gradients, preserving the constant Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$; and (ii) it is semantically natural, since $\pi_{\mathrm{old}}$ is already frozen as the reference policy in each iteration.
+Remark 4.4 (Two-Geometry Structure).
+With this derivation, the entire OPO framework admits a clean geometric decomposition:
+Alignment == $\alpha$-geometry (sampling) $\times\;$ mirror geometry (optimization)
+The sampling axis is an $\alpha$-divergence interpolation in the statistical manifold; the optimization axis is a Bregman divergence projection in ratio space.
+These two geometries are structurally orthogonal: $\alpha$ shapes the driving force without altering the Hessian, and $\mu$ shapes the curvature without altering the target.
 4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
@@
 $D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (8)
-This choice transforms the alignment problem into a quadratic optimization in ratio space.
-4.4 Orthogonalized Policy Optimization (OPO)
-Substituting the Euclidean geometry into our framework yields the OPO objective:
+This choice transforms the alignment problem into a quadratic optimization in ratio space, with curvature $\mu I$ independent of the data.
+4.4 The OPO Objective
+Combining the $\alpha$-escort sampling weight (Axis 1) with the Euclidean mirror map (Axis 2) yields the OPO objective:
 $\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (9)
-where $\mu=1/\eta$ is the stiffness parameter. The solution is available in closed form: $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
+where $\mu=1/\eta$ is the stiffness parameter, with closed-form solution $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
+4.5 Gradient Dynamics: Why OPO Does Not Saturate
+Differentiating Eq. (9) w.r.t. $v_{\theta}(y)$:
+$\nabla_{v}\mathcal{L}_{\text{OPO}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (10)
+with unique equilibrium $v^{*}(y)=\omega_{\alpha}(y)/\mu$. This linear relationship has three desirable properties:
+  1. 1.
+    Stable Equilibrium: Strict convexity in $v_{\theta}$ yields a unique stationary point.
+  2. 2.
+    No Saturation: Unlike sigmoid-based losses, the gradient does not vanish as $v$ grows—it increases proportionally to the distance from equilibrium.
+  3. 3.
+    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the curvature. The two do not interact.
+4.6 Implementation: Log-Ratio Approximation
+The ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$ is not directly available from LLM logits.
+However, for the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ (which is directly computable), the first-order Taylor expansion $v=e^{\Delta}-1\approx\Delta$ holds within the trust region.
+This yields the implementable form:
+$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (11)
+This retains the linear dynamics of Euclidean geometry while operating entirely in log-space.
+The approximation is justified by on-policy anchoring: setting $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ ensures $\Delta\approx 0$ at the start of each iteration.
+4.7 Constructive Derivation: Regression in Ratio Space
+We provide an alternative, constructive derivation that arrives at the same OPO loss from pure first principles—no mirror maps required.
+Step 1: Define the target offset.
+Given the escort weight $\omega_{\alpha}$ and regularization strength $\mu$, the optimal ratio offset is $v^{*}(y):=\omega_{\alpha}(y)/\mu$.
+This is the ideal deviation from the reference: the quality signal drives the direction; $\mu$ bounds the magnitude.
+Step 2: Minimize the distance.
+The natural metric in ratio space is mean squared error:
+$\mathcal{J}(\theta)=\frac{\mu}{2}\big(v_{\theta}(y)-v^{*}(y)\big)^{2}$ (12)
+Step 3: Complete the square.
+Expanding and discarding the $\theta$-independent constant $\omega_{\alpha}^{2}/(2\mu)$:
+$\mathcal{J}(\theta)=\underbrace{\frac{\mu}{2}\,v_{\theta}^{2}}_{\text{quadratic regularizer}}-\underbrace{\omega_{\alpha}\cdot v_{\theta}}_{\text{linear drive}}+\text{const}\;\;\Longrightarrow\;\;\mathcal{L}(\theta)=-\omega_{\alpha}\,v_{\theta}+\frac{\mu}{2}\,v_{\theta}^{2}$ (13)
+which is exactly the OPO objective (Eq.
+9).
+The completing-the-square identity provides a constructive proof: MSE regression in ratio space is gradient-equivalent to the linear-plus-quadratic OPO loss.
+Remark 4.5 (Equivalence of Perspectives).
+[…]
 Remark 4.6 (Analogy with Quantum Amplitudes).
-A physical analogy elucidates the advantages of the Euclidean geometry.
-  • Ratio Space as Amplitude: In quantum mechanics, superpositions are linear in amplitude ($\psi$).
-Similarly, OPO operates in the ratio space $v$, allowing preference signals to superpose linearly without saturation.
-  • KL Space as Probability: KL-based methods operate in log-probability space.
-Analogous to adding energies in a Boltzmann distribution, this enforces exponential scaling, which leads to "probability collapse" or saturation when signals are strong.
-By working in "amplitude" (ratio) space with quadratic "energy" (regularization), OPO maintains a non-saturating linear response field, crucial for long-chain reasoning where we must avoid premature collapse to a single path.
-4.5 Log-Ratio Approximation
-For implementation in LLMs (which output logits), we use the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$. As shown in Appendix A, for small trust regions, $v\approx\Delta$.
-$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\sum_{y\in S}\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (8)
-This retains the linear dynamics of the Euclidean geometry while remaining compatible with standard gradient-based training.
-Remark 4.2 (On-Policy Anchoring).
-We set $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ (on-policy anchoring), ensuring the trust region condition $\Delta\approx 0$ holds locally, justifying the approximation.
-4.6 Gradient Dynamics
-For the ratio-coordinate objective Equation˜7, the gradient w.r.t. $v_{\theta}(y)$ is:
-$\nabla_{v}\mathcal{L}_{\text{OPO}}^{\text{ratio}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (9)
-with the unique equilibrium:
-$v^{*}(y)=\frac{\omega_{\alpha}(y)}{\mu}$ (10)
-This linear relationship has several desirable properties:
-  1. 1.
-    Stable Equilibrium: The objective is strictly convex in $v_{\theta}(y)$ for fixed $\omega_{\alpha}(y)$, yielding a unique stationary point in the ratio coordinate.
-  2. 2.
-    No Saturation: Unlike sigmoid-based losses, the gradient magnitude does not vanish as $v$ grows.
-  3. 3.
-    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the regularization strength. The two do not interact.
+A physical analogy: in quantum mechanics, superpositions are linear in amplitude ($\psi$), while measurements are quadratic in amplitude.
+Similarly, OPO operates in the ratio space $v$ (“amplitude”), where preference signals superpose linearly, and regularization is quadratic.
+KL-based methods operate in log-probability (“energy”) space, where strong signals lead to exponential collapse.
+The amplitude picture explains why OPO sustains non-saturating gradients in high-confidence regimes.
 5 Theoretical Analysis
-We formlize the decoupling of design axes and analyze the distinct regularization properties of OPO.
-5.1 Decoupled Sampling and Optimization Geometry
-We analyze the optimization dynamics in the function space of ratio outputs $v$.
-This perspective captures the intrinsic objective geometry.
-We first formalize the connection to parameter space dynamics.
+The preceding section derived OPO from two complementary perspectives (mirror descent and ratio regression).
+We now prove that OPO’s key structural properties—orthogonal control and global contraction—hold rigorously.
+5.1 Formal Decoupling of Geometry Axes
+All analysis is conducted in the function space of ratio outputs $v\in L^{2}(\pi_{\mathrm{ref}})$, which captures the intrinsic objective geometry. We first bridge to parameter space.
 Lemma 5.1 (Parameter Space Gradient Projection).
@@
 Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
-This structural decoupling guarantees stable, predictable convergence dynamics in the function space.
+Remark 5.3 (Scope of Orthogonality).
+The structural orthogonality in Theorem 5.2 holds exactly in the ratio function space $L^{2}(\pi_{\mathrm{ref}})$.
+In parameter space $\theta$, interaction may arise through policy parametrization, since $\nabla_{\theta}v=\nabla_{\theta}(\exp(\Delta)-1)$ depends on the training trajectory influenced by $\alpha$.
+However, this coupling is a property of the parametrization, not the intrinsic objective curvature: the Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$ remains structurally independent of $\alpha$ regardless of the parameter mapping.
+A natural concern is whether the stop-gradient operator in $\omega_{\alpha}$ (Remark 4.3) introduces bias. The following lemma shows it does not, at the anchor point:
+Lemma 5.4 (Exactness of Surrogate Gradient at Anchor Point).
+Let $\mathcal{L}_{\mathrm{true}}(\theta)$ be the exact objective without stop-gradient on $\omega_{\alpha}$, and $\mathcal{L}_{\mathrm{OPO}}(\theta)$ the surrogate with $\operatorname{sg}$.
+Under on-policy anchoring $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, at the start of each iteration ($\theta=\theta_{\mathrm{old}}$):
+$\left.\nabla_{\theta}\mathcal{L}_{\mathrm{OPO}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}=\left.\nabla_{\theta}\mathcal{L}_{\mathrm{true}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}$ (17)
+Furthermore, the gradient error grows only as $\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$.
+Proof.
+By the product rule, the true function-space gradient contains a cross-term $(\nabla_{v}\omega_{\alpha})\,v$ that the surrogate omits.
+At the anchor point $\pi_{\theta}=\pi_{\mathrm{old}}$, the ratio offset is exactly $v=\pi_{\theta}/\pi_{\mathrm{old}}-1=0$, so this cross-term vanishes.
+For $\theta\neq\theta_{\mathrm{old}}$, the error is bounded by $\|(\nabla_{v}\omega_{\alpha})\,v\|=\mathcal{O}(\|v\|)=\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$, which is small within the trust region.
+∎
+This confirms that stop-gradient is not an engineering hack but a first-order exact approximation: the surrogate and true objectives agree to first order at every on-policy anchor.
+Under sufficiently small trust-region updates (i.e., $\|v\|$ bounded by the $\chi^{2}$ penalty), the cumulative bias across iterations remains second-order small, since each anchor reset re-zeroes the first-order error term.
 Corollary 5.5 (Global Contraction and Linear Convergence).
-Gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
+Under the assumption that the policy class can represent $v^{*}$ and that local linearization of the map $\theta\mapsto v_{\theta}$ holds near the optimum, gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
 $v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (18)
@@
 $|\nabla_{\text{OPO}}|\geq\mu\delta$ (21)
-This guarantees a linear driving force that does not vanish arbitrarily, preventing the saturation observed in KL-based methods.
+This guarantees a linear driving force that does not vanish arbitrarily.
+The practical consequence is stark: in mathematical reasoning, models frequently assign near-1.0 probability to correct intermediate steps.
+KL-based methods penalize these high-confidence correct outputs by exponentially shrinking the learning signal, whereas OPO maintains a constant reward-driven pull via its linear dynamics—allowing the model to continue refining even its most confident predictions.
 5.4 Algorithm and Flowchart
@@
 Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1.
-Ratio $r=\pi/\pi_{old}$ 2.
-Reweight $\omega_{\alpha}\leftarrow r^{\alpha}A$ 3.
+Ratio $v=\pi/\pi_{old}-1$ 2.
+$\alpha$-Weight $\omega_{\alpha}=(e^{A}/\tilde{p})^{1-\alpha}$ 3.
 Orth.
-Penalty $\mu(r-1)^{2}$ Update $\theta$ Next Iteration
+Penalty $\mu\,v^{2}$ Update $\theta$ Next Iteration
 Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework.
-The core innovation lies in the “OPO Decomposition” block (green), which separates the advantage signal (reweighted by $\alpha$) from the trust-region constraint (quadratic penalty $\chi^{2}$), ensuring linear non-saturating gradients.
-Algorithm 1: Orthogonalized Policy Optimization (OPO)  Input: Initial policy $\pi_{\theta}$, hyperparameters $\alpha,\mu,\eta$ for each iteration do 1.
-Set Reference: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ (on-policy anchoring) 2.
-Collect Rollouts: Sample $(x,y)\sim\pi_{\theta}$, compute rewards $R(x,y)$ 3.
-Compute Advantages: $A(y)\leftarrow R(y)-\bar{R}$ (group-normalized) 4.
-Compute Weights: $\omega_{\alpha}(y)\leftarrow A(y)\cdot[\text{reweight by }\alpha]$ (sampling geometry) 5.
-Compute Ratio: $v_{\theta}(y)\leftarrow\pi_{\theta}(y)/\pi_{\mathrm{ref}}(y)-1$ (or use log-approx) 6.
-Compute Loss: $\mathcal{L}\leftarrow-\sum_{y}\omega_{\alpha}(y)v_{\theta}(y)+\frac{\mu}{2}\sum_{y}v_{\theta}(y)^{2}$ (optimization geometry) 7.
-Update Policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$ end for
-Figure 2: Pseudocode for Orthogonalized Policy Optimization. The algorithm explicitly decouples sampling geometry (step 4) from optimization geometry (step 6).
+The “OPO Decomposition” block (green) embodies the two-geometry structure: the $\alpha$-divergence projection (step 2) governs sampling, while the $\chi^{2}$ penalty (step 3) governs optimization.
+These two axes are structurally orthogonal.
+Algorithm 1 Orthogonalized Policy Optimization (OPO)
+1:Initial policy $\pi_{\theta}$; parameters $\alpha\in[0,1],\;\mu>0,\;\eta>0$
+2:for each iteration do
+3:  Anchor: $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$ $\triangleright$ On-policy anchoring
+4:  Rollout: Sample $(x,y)\sim\pi_{\theta}$; compute rewards $R(x,y)$
+5:  Advantage: $A(y)\leftarrow\mathrm{Normalize}(R(x,y))$ $\triangleright$ Group-normalized
+6:  Escort Weight: $\omega_{\alpha}(y)\leftarrow\exp\!\big((1{-}\alpha)\,A(y)\big)\cdot\operatorname{sg}\!\big(\pi_{\mathrm{ref}}(y)^{-(1-\alpha)}\big)$ $\triangleright$ Eq. 6
+7:  Log-Ratio: $\Delta_{\theta}(y)\leftarrow\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$
+8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$ $\triangleright$ $y$: complete response sequence
+9:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
+10:end for
 6 Experiments
@@
       Overlong Reward Shaping: Applies a length-aware penalty and masks gradients for truncated responses to reduce noise.
-  • OPO: The proposed method with $\alpha=0.6$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$.
+  • OPO: The proposed method with $\alpha=0.4$, $\mu=1.0$, on-policy anchoring ($\pi_{\mathrm{ref}}=\pi_{\text{old}}$), and adaptive $\tau$ with range $[0.2,1.5]$.
+Note: $1-\alpha=0.6$ controls the effective advantage amplification in Eq.
+(6).
 6.2 Results
@@
 Table 1 summarizes the key metrics across all four algorithms.
-Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3. Reward and gradient norm are averaged over the final 20 training steps.
+Table 1: Comparison of alignment algorithms on Qwen3-1.7B + MATH Level 3.
+Reward and gradient norm are averaged over the final 20 training steps.
+Higher gradient norm at convergence, combined with stable (non-exploding) training dynamics, indicates absence of gradient saturation rather than optimization instability.
 Algorithm Mean Reward Grad Norm Characteristics
@@
 Figure 3 illustrates the gradient norm dynamics across training.
-OPO exhibits consistently higher gradient norms, validating the non-saturating property of the $\chi^{2}$ geometry.
+OPO exhibits consistently higher but stable (non-exploding) gradient norms throughout training, validating the non-saturating property of the $\chi^{2}$ geometry.
+The stability of these norms rules out the alternative explanation of optimization instability or under-regularization.
 DAPO shows severely diminished gradients, explaining its early plateau.
@@
 We attribute this to the granularity of supervision: GRPO and DAPO utilize token-level updates, imposing dense supervision that forces the policy to collapse rapidly into specific phrasings.
-In contrast, OPO (and GSPO) employs sequence-level updates.
+In contrast, OPO (and GSPO) employs sequence-level updates, where $y$ denotes an entire response trajectory and the loss is averaged over complete sequences rather than individual tokens.
 This “sparse” supervision aligns the total trajectory probability without micromanaging individual tokens, thereby preserving policy diversity and exploration potential.
@@
 Relation to Existing Methods.
-OPO can be viewed as a generalization of existing alignment methods:
-  • Setting $\alpha=1$ and removing the quadratic term recovers SFT-like behavior.
-  • Setting $\alpha\to 0$ and using KL geometry yields objectives with similar peak-seeking and exponential-type weighting effects as in KL-regularized preference optimization (e.g., DPO).
-  • OPO occupies the previously unexplored region of the design space: peak-seeking sampling with quadratic regularization in ratio coordinates.
+OPO can be viewed as occupying a previously unexplored point in the design space:
+  • $\alpha\to 1$ (conservative escort): $\omega_{\alpha}\to 1$, recovering supervised fine-tuning–like uniform weighting.
+  • $\alpha\to 0$ (aggressive escort) with KL geometry: the exponential-type weighting resembles KL-regularized preference optimization (e.g., DPO), but inherits gradient saturation.
+  • OPO: moderate $\alpha$ (peak-seeking escort) combined with Euclidean ($\chi^{2}$) geometry—non-saturating gradients with advantage-aware sampling.
 Computational Cost.
@@
 Implementation Notes.
-The OPO objective (Equation˜7) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
+The OPO objective (Equation˜9) involves expectations under $\pi_{\mathrm{ref}}$. In practice:
   • For on-policy training where samples come from $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, the batch itself provides an unbiased estimate.
@@
 8 Conclusion
-We have presented Orthogonalized Policy Optimization (OPO), a framework that decomposes alignment objectives into two independent design axes: sampling geometry and optimization geometry.
-By working in ratio coordinates where Pearson $\chi^{2}$ divergence naturally induces quadratic penalties, OPO avoids the entangled dynamics of KL-based methods and provides a well-conditioned objective with linear gradient dynamics.
-Our analysis shows that OPO maintains stable optimization even in high-confidence regimes, addressing a key limitation of existing methods in reasoning tasks.
+We have presented Orthogonalized Policy Optimization (OPO), a framework grounded in two independent geometries: an $\alpha$-divergence escort projection that shapes sampling, and a Euclidean mirror map (Pearson $\chi^{2}$) that shapes optimization.
+The same loss emerges from two complementary derivations—mirror descent and ratio-space regression—confirming it is uniquely determined by the geometry choice.
+The resulting objective has a constant Hessian, linear non-saturating gradients, and global contraction, addressing a key limitation of KL-based methods in high-confidence reasoning tasks.
 We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
@@
 Let $\delta=\|\Delta_{\theta}\|_{\infty}$.
-For $\delta<1$ (trust region regime):
+For sufficiently small $\delta$ (in particular $\delta<1$, which is enforced by the $\chi^{2}$ trust-region penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ bounding $\|\Delta\|_{\infty}$):
 $|v_{\theta}-\Delta_{\theta}|\leq\frac{1}{2}\delta^{2}e^{\delta},\quad|v_{\theta}^{2}-\Delta_{\theta}^{2}|\leq\delta^{3}e^{2\delta}$ (22)
@@
 This duality implies that $\alpha$ purely shapes the objective (alignment direction), while $\mu$ purely enforces the feasibility radius (optimization geometry).
-A.3 Distributional Stability Guarantees
+A.3 Proof of Proposition 4.2: $\alpha$-Geometric Interpolation
+Proof.
+Define the oracle target $P^{*}(y)\propto\exp(A(y))$ and the parameterized family $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$ for $\alpha\in[0,1]$.
+Step 1 (Geodesic structure). In the $e$-affine coordinate system of the statistical manifold [11], the log-density of $\rho_{\alpha}$ is:
+$\log\rho_{\alpha}(y)=\alpha\log\tilde{p}(y)+(1-\alpha)\log P^{*}(y)-\log Z_{\alpha}$ (24)
+where $Z_{\alpha}$ is the normalizing constant.
+Since $\log\rho_{\alpha}$ is an affine function of $\alpha$ in the natural parameter, $\{\rho_{\alpha}\}_{\alpha\in[0,1]}$ forms an $e$-geodesic (exponential geodesic) connecting $P^{*}$ (at $\alpha=0$) to $\tilde{p}$ (at $\alpha=1$) in the exponential family.
+Step 2 (Importance weight). The sampling weight is:
+$\omega_{\alpha}(y)=\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\propto\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}$ (25)
+In the on-policy setting where $P^{*}(y)=\exp(A(y))$ and $\tilde{p}=\pi_{\mathrm{old}}$, this yields $\omega_{\alpha}\propto(\exp(A)/\tilde{p})^{1-\alpha}$, which simplifies to $\exp((1-\alpha)A)$ when the $\tilde{p}$ correction is absorbed by the sampling measure.
+Step 3 (Geometric consistency).
+This interpolation corresponds to the exponential geodesic in the statistical manifold and is consistent with Rényi $\alpha$-divergence interpolation geometry [11].
+Under the dually flat structure of exponential families, it coincides with the $e$-projection onto the geodesic submanifold connecting $\tilde{p}$ and $P^{*}$.
+∎
+A.4 Distributional Stability Guarantees
 Proposition A.3 ($\chi^{2}$ Controls Total Variation).

### v4

@@
 Abstract
-Large Language Model (LLM) alignment objectives are often presented as a collection of distinct algorithms—PPO, DPO, IPO, and their variants—each motivated by different derivations.
-In this work, we argue that this diversity obscures a simpler underlying structure.
-At a fundamental level, alignment objectives involve two independent design choices: (i) how training signals are sampled and weighted, and (ii) how deviations from a reference policy are geometrically penalized.
-Existing methods typically entangle these choices through a single divergence, most commonly the Kullback–Leibler (KL) divergence.
-We show that this entanglement is not merely a modeling convenience but a source of systematic instability.
-When the same divergence simultaneously determines sample weighting and optimization curvature, adjusting one aspect (e.g., exploration strength) inevitably alters the other (e.g., gradient geometry).
-This coupling is particularly problematic in preference-based reinforcement learning, where advantage signals are unbounded and high-confidence regimes are common.
-We propose a principled structural remedy: formulate alignment as an orthogonal mirror descent problem, in which sampling geometry enters as a linear driving force derived from an $\alpha$-divergence projection, while optimization geometry is determined independently by a Bregman divergence (mirror map).
-This perspective leads to a new alignment objective—Orthogonalized Policy Optimization (OPO)—obtained by choosing a Euclidean mirror map in likelihood-ratio space.
-The resulting objective admits a closed-form solution, linear (non-saturating) gradient dynamics, and a well-conditioned trust region, while remaining fully compatible with standard LLM training pipelines.
+We present Orthogonalized Policy Optimization (OPO), a unified theoretical account of LLM alignment grounded in a work–dissipation principle.
+The policy update is characterized as a constrained proximal response that maximizes external work induced by an $\alpha$-escort sampling field, while paying an intrinsic dissipation cost given by a quadratic fluctuation energy in $\chi^{2}$ (ratio) geometry.
+This single variational principle simultaneously admits three equivalent interpretations: (i) a mirror-descent step with a Euclidean mirror map in ratio space, (ii) a Hilbert-space projection via the orthogonal projection theorem in $L^{2}(\pi_{k})$, and (iii) a linear-response law from near-equilibrium statistical mechanics.
+Their convergence to the same closed-form update confirms that OPO is the unique quadratic proximal response within ratio geometry.
+The framework cleanly decouples sampling geometry ($\alpha$) from optimization geometry ($\mu$), yields constant Hessian and non-saturating linear gradients, and reveals that advantage $z$-score normalization is not a heuristic but a conservation-law projection.
+Experiments on mathematical reasoning tasks demonstrate that OPO outperforms GRPO, GSPO, and DAPO while maintaining healthy gradient dynamics throughout training.
 1 Introduction
@@
 This manifests as gradient saturation, where the driving force for further improvement vanishes exponentially as the model becomes confident.
-Consequently, training often plateaus prematurely, behavior commonly attributed to "over-regularization" but which we argue is intrinsic to the chosen geometry.
+Consequently, training often plateaus prematurely, behavior commonly attributed to “over-regularization” but which we argue is intrinsic to the chosen geometry.
 We contend that this issue stems from a structural conflation in existing objective designs.
@@
 Adjusting one inevitably perturbs the other, creating a dilemma where aggressive sampling destabilizes training, while stable optimization stifles exploration.
-To resolve this, we propose a unifying framework based on Orthogonal Mirror Descent.
-By viewing alignment as a mirror descent process, we can explicitly decouple these two axes: the sampling geometry acts as a linear driving force defined by an advantage-weighted distribution, while the optimization geometry is independently governed by a Bregman divergence (mirror map).
-This perspective reveals that the saturation observed in DPO and PPO is not a necessary feature of alignment, but a specific consequence of using the Negative Entropy mirror map (which induces KL divergence).
-Guided by this framework, we introduce Orthogonalized Policy Optimization (OPO), a new alignment objective derived by selecting the Euclidean mirror map in the space of likelihood ratios.
-Unlike the exponential geometry of KL, the Euclidean geometry induces a Pearson $\chi^{2}$ trust region, leading to a quadratic regularizer that maintains a linear, non-saturating driving force even in high-confidence regimes.
-This formulation cleanly separates the sampling intensity (controlled by an $\alpha$-divergence parameter) from the regularization strength (controlled by a stiffness parameter), allowing for robust, continuous improvement.
+To resolve this, we propose a unifying framework based on a work–dissipation principle: the policy update maximizes external work induced by an $\alpha$-escort sampling field, while paying an intrinsic dissipation cost in $\chi^{2}$ ratio geometry.
+This single variational principle cleanly decouples the two axes and admits three equivalent mathematical interpretations—mirror descent, Hilbert-space projection, and linear response—each illuminating a different structural facet.
 Our contributions are as follows:
   • We identify the implicit coupling of sampling and optimization geometries in KL-based alignment methods as a root cause of gradient saturation.
-  • We propose the Orthogonal Mirror Descent framework, which decouples these design axes via two independent geometries: an $\alpha$-divergence projection governs sampling, while a Bregman divergence (mirror map) governs optimization.
-  • We derive Orthogonalized Policy Optimization (OPO), a simple, closed-form objective based on Euclidean geometry that ensures linear gradient dynamics and global contraction.
-  • We empirically demonstrate that OPO outperforms strong baselines like DPO and GRPO on mathematical reasoning tasks, validating the benefits of non-saturating gradients.
+  • We propose a unified work–dissipation principle that decouples these design axes: $\alpha$-escort shapes the driving force, while a $\chi^{2}$ quadratic penalty shapes the optimization curvature.
+  • We show that the resulting OPO objective admits three equivalent interpretations—Bregman mirror descent, Hilbert-space orthogonal projection, and near-equilibrium linear response—confirming it is the unique quadratic proximal response in ratio geometry.
+The Hilbert derivation further reveals that advantage $z$-score normalization is a conservation-law projection.
+  • We empirically demonstrate that OPO outperforms strong baselines (GRPO, GSPO, DAPO) on mathematical reasoning tasks.
+A gradient response efficiency analysis—normalizing gradient magnitude by the advantage signal amplitude—reveals that OPO maintains ${\sim}5{\times}$ higher gradient-per-signal than KL-based baselines, providing direct quantitative evidence for the linear response prediction.
 2 Related Work
@@
 OPO extends this by replacing the KL-based geometry with a quadratic ($\chi^{2}$) geometry in ratio coordinates, providing a different form of implicit regularization.
-3 Preliminaries: Two Coordinate Systems
-Before presenting the OPO framework, we establish the two coordinate systems that underlie all ratio-based alignment objectives.
-3.1 Ratio Space vs. Log-Ratio Space
-Let the reference policy be $\pi_{\mathrm{ref}}$. We define two coordinate systems for measuring policy deviation:
-Definition 3.1 (Ratio Coordinates).
-The density ratio and centered ratio are:
-$t_{\theta}(y):=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)},\qquad v_{\theta}(y):=t_{\theta}(y)-1=\frac{\pi_{\theta}(y)}{\pi_{\mathrm{ref}}(y)}-1$ (1)
-Definition 3.2 (Log-Ratio Coordinates).
-The log-ratio (natural parameterization of LLMs) is:
-$\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$ (2)
-so that $t_{\theta}(y)=\exp(\Delta_{\theta}(y))$ and $v_{\theta}(y)=\exp(\Delta_{\theta}(y))-1$.
-The key observation is that $\chi^{2}$ divergence is naturally quadratic in ratio coordinates $v$, while KL divergence involves the log-ratio $\Delta$.
-This distinction is crucial for understanding the gradient dynamics of different alignment objectives.
-4 Theoretical Framework: Orthogonal Mirror Descent
-We present a unified framework that fundamentally decouples the preference sampling geometry (what to learn) from the optimization geometry (how to learn).
-We formulate policy alignment as a functional Orthogonal Mirror Descent problem.
-4.1 The General Framework
-We view alignment as an iterative projection process.
-At each step $k$, we seek a policy $\pi$ that maximizes the expected return under a specific sampling distribution $\rho_{\alpha}$, subject to a trust-region constraint defined by a divergence $D_{\Psi}$:
-$\pi_{k+1}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{y\sim\rho_{\alpha}}\left[A(y)\log\pi(y)\right]-\frac{1}{\eta}D_{\Psi}(\pi\|\pi_{k})$ (3)
-This formulation cleanly orthogonizes two design axes:
-  1. 1.
-    Sampling Geometry ($\rho_{\alpha}$): Determines the effective weighting of samples (the linear driving force).
-  2. 2.
-    Optimization Geometry ($D_{\Psi}$): Determines the mirror map (regularization/curvature) used to project the update.
-4.2 Axis 1: Sampling Geometry via $\alpha$-Divergence Projection
-The sampling weight $\omega_{\alpha}$ determines the effective weighting of training examples.
-We derive it from a principled $\alpha$-divergence interpolation, providing the sampling axis with the same geometric rigor as the optimization axis.
-Definition 4.1 (Oracle Target Distribution).
-Given a non-negative quality signal $Q(y)\geq 0$ and the behavior policy $\tilde{p}(y):=\pi_{\mathrm{old}}(y)$, define the oracle target distribution:
-$P^{*}(y)\;\propto\;Q(y)$ (4)
-In the standard RL setting with advantage function $A(y)$, the natural choice is $Q(y)=\exp(A(y))$, corresponding to the Boltzmann-optimal density.
-The importance sampling (IS) ratio is $w_{\mathrm{IS}}(y)=P^{*}(y)/\tilde{p}(y)\propto Q(y)/\tilde{p}(y)$.
-Direct use of $w_{\mathrm{IS}}$ suffers from variance explosion in large action spaces.
-The parameter $\alpha\in[0,1]$ geometrically tempers this correction by interpolating along the $\alpha$-geodesic [11] connecting $\tilde{p}$ to $P^{*}$:
-Proposition 4.2 ($\alpha$-Escort Weighting).
-The $\alpha$-geodesic between $\tilde{p}$ and $P^{*}$ yields the escort distribution $\rho_{\alpha}(y)\propto\tilde{p}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}$. The corresponding importance weight is:
-$\boxed{\omega_{\alpha}(y)\;=\;\frac{\rho_{\alpha}(y)}{\tilde{p}(y)}\;\propto\;\left(\frac{P^{*}(y)}{\tilde{p}(y)}\right)^{1-\alpha}=\left(\frac{\exp(A(y))}{\tilde{p}(y)}\right)^{1-\alpha}}$ (5)
-This parameterization has clear geometric semantics:
-  • $\alpha\to 1$: $\omega_{\alpha}\to 1$ (uniform weighting; mode-covering, conservative).
-  • $\alpha\to 0$: $\omega_{\alpha}\to P^{*}/\tilde{p}$ (full IS correction; peak-seeking, aggressive).
-  • $0<\alpha<1$: Smooth interpolation governed by $\alpha$-geometry.
-Remark 4.3 (On-Policy Implementation with Frozen Denominator).
-In the on-policy setting where $\tilde{p}=\pi_{\mathrm{old}}$, the theoretical escort weight $\omega_{\alpha}=(\exp(A)/\tilde{p})^{1-\alpha}$ can be decomposed as:
-$\omega_{\alpha}(y)\;=\;\exp\!\big((1-\alpha)\,A(y)\big)\;\cdot\;\operatorname{sg}\!\left(\frac{1}{\pi_{\mathrm{old}}(y)^{\,1-\alpha}}\right)$ (6)
-where $\operatorname{sg}(\cdot)$ denotes the stop-gradient operator.
-This formulation is not an approximation—it computes the exact escort weight from Eq.
-(5).
-The stop-gradient on the $1/\pi_{\mathrm{old}}^{1-\alpha}$ factor serves two critical purposes: (i) it prevents the $1/p$ denominator from injecting unbounded gradients, preserving the constant Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$; and (ii) it is semantically natural, since $\pi_{\mathrm{old}}$ is already frozen as the reference policy in each iteration.
-Remark 4.4 (Two-Geometry Structure).
-With this derivation, the entire OPO framework admits a clean geometric decomposition:
-Alignment == $\alpha$-geometry (sampling) $\times\;$ mirror geometry (optimization)
-The sampling axis is an $\alpha$-divergence interpolation in the statistical manifold; the optimization axis is a Bregman divergence projection in ratio space.
-These two geometries are structurally orthogonal: $\alpha$ shapes the driving force without altering the Hessian, and $\mu$ shapes the curvature without altering the target.
-4.3 Axis 2: Optimization Geometry: Euclidean Mirror Map
-The core innovation of OPO is the choice of the mirror map.
-Instead of the standard Negative Entropy map (which leads to KL divergence and exponential updates), we choose the Euclidean Mirror Map in the space of likelihood ratios $v=\pi/\pi_{\mathrm{ref}}-1$:
-$\Psi(v)=\frac{1}{2}\|v\|_{L^{2}(\pi_{\mathrm{ref}})}^{2}=\frac{1}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}[v(y)^{2}]$ (7)
-The induced Bregman divergence is the Pearson $\chi^{2}$ divergence:
-$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\left[\left(\frac{\pi}{\pi_{\mathrm{ref}}}-\frac{\pi_{k}}{\pi_{\mathrm{ref}}}\right)^{2}\right]=D_{\chi^{2}}(\pi\|\pi_{k})$ (8)
-This choice transforms the alignment problem into a quadratic optimization in ratio space, with curvature $\mu I$ independent of the data.
-4.4 The OPO Objective
-Combining the $\alpha$-escort sampling weight (Axis 1) with the Euclidean mirror map (Axis 2) yields the OPO objective:
-$\boxed{\mathcal{L}_{\text{OPO}}=-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\omega_{\alpha}(y)\,v(y)\right]+\frac{\mu}{2}\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[v(y)^{2}\right]}$ (9)
-where $\mu=1/\eta$ is the stiffness parameter, with closed-form solution $v^{*}(y)=\omega_{\alpha}(y)/\mu$.
-4.5 Gradient Dynamics: Why OPO Does Not Saturate
-Differentiating Eq. (9) w.r.t. $v_{\theta}(y)$:
-$\nabla_{v}\mathcal{L}_{\text{OPO}}=-\omega_{\alpha}(y)+\mu\,v_{\theta}(y)$ (10)
-with unique equilibrium $v^{*}(y)=\omega_{\alpha}(y)/\mu$. This linear relationship has three desirable properties:
-  1. 1.
-    Stable Equilibrium: Strict convexity in $v_{\theta}$ yields a unique stationary point.
-  2. 2.
-    No Saturation: Unlike sigmoid-based losses, the gradient does not vanish as $v$ grows—it increases proportionally to the distance from equilibrium.
-  3. 3.
-    Orthogonal Control: Changing $\alpha$ (sampling) affects only $\omega_{\alpha}$; changing $\mu$ (optimization) affects only the curvature. The two do not interact.
-4.6 Implementation: Log-Ratio Approximation
-The ratio $v_{\theta}=\pi_{\theta}/\pi_{\mathrm{ref}}-1$ is not directly available from LLM logits.
-However, for the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ (which is directly computable), the first-order Taylor expansion $v=e^{\Delta}-1\approx\Delta$ holds within the trust region.
-This yields the implementable form:
-$\boxed{\mathcal{L}_{\text{OPO}}^{\text{log}}\approx-\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}\big[\Delta_{\theta}(y)^{2}\big]}$ (11)
-This retains the linear dynamics of Euclidean geometry while operating entirely in log-space.
-The approximation is justified by on-policy anchoring: setting $\pi_{\mathrm{ref}}=\pi_{\text{old}}$ ensures $\Delta\approx 0$ at the start of each iteration.
-4.7 Constructive Derivation: Regression in Ratio Space
-We provide an alternative, constructive derivation that arrives at the same OPO loss from pure first principles—no mirror maps required.
-Step 1: Define the target offset.
-Given the escort weight $\omega_{\alpha}$ and regularization strength $\mu$, the optimal ratio offset is $v^{*}(y):=\omega_{\alpha}(y)/\mu$.
-This is the ideal deviation from the reference: the quality signal drives the direction; $\mu$ bounds the magnitude.
-Step 2: Minimize the distance.
-The natural metric in ratio space is mean squared error:
-$\mathcal{J}(\theta)=\frac{\mu}{2}\big(v_{\theta}(y)-v^{*}(y)\big)^{2}$ (12)
-Step 3: Complete the square.
-Expanding and discarding the $\theta$-independent constant $\omega_{\alpha}^{2}/(2\mu)$:
-$\mathcal{J}(\theta)=\underbrace{\frac{\mu}{2}\,v_{\theta}^{2}}_{\text{quadratic regularizer}}-\underbrace{\omega_{\alpha}\cdot v_{\theta}}_{\text{linear drive}}+\text{const}\;\;\Longrightarrow\;\;\mathcal{L}(\theta)=-\omega_{\alpha}\,v_{\theta}+\frac{\mu}{2}\,v_{\theta}^{2}$ (13)
-which is exactly the OPO objective (Eq.
-9).
-The completing-the-square identity provides a constructive proof: MSE regression in ratio space is gradient-equivalent to the linear-plus-quadratic OPO loss.
-Remark 4.5 (Equivalence of Perspectives).
-[…]
-Remark 4.6 (Analogy with Quantum Amplitudes).
-A physical analogy: in quantum mechanics, superpositions are linear in amplitude ($\psi$), while measurements are quadratic in amplitude.
-Similarly, OPO operates in the ratio space $v$ (“amplitude”), where preference signals superpose linearly, and regularization is quadratic.
-KL-based methods operate in log-probability (“energy”) space, where strong signals lead to exponential collapse.
-The amplitude picture explains why OPO sustains non-saturating gradients in high-confidence regimes.
-5 Theoretical Analysis
-The preceding section derived OPO from two complementary perspectives (mirror descent and ratio regression).
-We now prove that OPO’s key structural properties—orthogonal control and global contraction—hold rigorously.
-5.1 Formal Decoupling of Geometry Axes
-All analysis is conducted in the function space of ratio outputs $v\in L^{2}(\pi_{\mathrm{ref}})$, which captures the intrinsic objective geometry. We first bridge to parameter space.
-Lemma 5.1 (Parameter Space Gradient Projection).
-Let $\mathcal{L}(v_{\theta})$ be a functional of the policy output $v_{\theta}(y)$.
-The gradient in parameter space $\theta$ is the projection of the function-space gradient $\nabla_{v}\mathcal{L}$ onto the tangent space of the policy manifold:
-$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{y\sim\pi_{\mathrm{ref}}}\left[\nabla_{v}\mathcal{L}(v_{\theta}(y))\cdot\nabla_{\theta}v_{\theta}(y)\right]$ (14)
-Thus, the geometric properties of $\nabla_{v}\mathcal{L}$ directly determine the driving force in parameter space, modulated by the policy’s local sensitivity $\nabla_{\theta}v_{\theta}$.
-With this link established, we characterize the geometry of $\mathcal{L}(v)$.
-Theorem 5.2 (Decoupled Sampling and Optimization Geometry).
-Consider the OPO objective $\mathcal{L}(v)=-\sum_{y}\omega_{\alpha}(y)v(y)+\frac{\mu}{2}\mathbb{E}_{\pi_{\mathrm{ref}}}[v(y)^{2}]$. In the space of ratio functions $v\in L^{2}(\pi_{\mathrm{ref}})$:
-  1. 1.
-    The first-order driving force depends on the sampling parameter $\alpha$:
-    $\nabla_{v}\mathcal{L}=-\omega_{\alpha}(y)+\mu v(y)$ (15)
-  2. 2.
-    The second-order curvature is independent of $\alpha$:
-    $\nabla^{2}_{v}\mathcal{L}=\mu I$ (16)
+3 Preliminaries: Ratio Coordinates and Conservation
+Let $\pi_{k}$ denote the anchor (reference) policy at iteration $k$, with $\pi_{k}(y)>0$ on its support.
+Definition 3.1 (Ratio and Log-Ratio Coordinates).
+Define
+$t(y):=\frac{\pi(y)}{\pi_{k}(y)},\qquad v(y):=t(y)-1=\frac{\pi(y)}{\pi_{k}(y)}-1,\qquad\Delta(y):=\log\pi(y)-\log\pi_{k}(y).$ (1)
+Then $t(y)=e^{\Delta(y)}$ and $v(y)=e^{\Delta(y)}-1$.
+Conservation law.
+Probability normalization $\sum_{y}\pi(y)=1$ is equivalent to
+$\mathbb{E}_{\pi_{k}}[t]=1\quad\Longleftrightarrow\quad\mathbb{E}_{\pi_{k}}[v]=0.$ (2)
+Hence admissible fluctuations lie in the closed subspace
+$\mathcal{H}_{0}:=\{v\in L^{2}(\pi_{k}):\mathbb{E}_{\pi_{k}}[v]=0\}.$ (3)
+4 Unified Work–Dissipation Principle
+This section provides the unifying core: OPO is a constrained proximal response that maximizes external work induced by sampling geometry, while paying an intrinsic dissipation (fluctuation) energy that stabilizes the update.
+4.1 Sampling Geometry as an $\alpha$-Escort Field
+Let $P^{*}$ be an oracle target distribution encoding preference/quality. A canonical choice is $P^{*}(y)\propto\exp(A(y))$ where $A$ is an advantage-like signal.
+Definition 4.1 ($\alpha$-Escort Distribution and Escort Weight).
+For $\alpha\in[0,1]$, define the escort distribution
+$\rho_{\alpha}(y)\propto\pi_{k}(y)^{\alpha}\,P^{*}(y)^{1-\alpha}.$ (4)
+Define the escort weight (Radon–Nikodym derivative)
+$\omega_{\alpha}(y):=\frac{\rho_{\alpha}(y)}{\pi_{k}(y)}.$ (5)
+Since $\rho_{\alpha}$ is normalized,
+$\mathbb{E}_{\pi_{k}}[\omega_{\alpha}]=1.$ (6)
+Proposition 4.2 ($\alpha$-Geometric Interpolation).
+The family $\{\rho_{\alpha}\}_{\alpha\in[0,1]}$ forms an $e$-geodesic connecting $P^{*}$ (at $\alpha=0$) to $\pi_{k}$ (at $\alpha=1$).
+The escort weight simplifies to $\omega_{\alpha}(y)\propto\bigl(\exp(A(y))/\pi_{k}(y)\bigr)^{1-\alpha}$.
+Geometric semantics.
+$\alpha$ controls how aggressively sampling emphasizes high-quality regions: $\alpha\to 1$ gives $\omega_{\alpha}\to 1$ (conservative); $\alpha\to 0$ gives $\omega_{\alpha}\to P^{*}/\pi_{k}$ (aggressive).
+On-policy implementation (stop-gradient).
+When $\pi_{k}=\pi_{\mathrm{old}}$ is frozen during an iteration,
+$\omega_{\alpha}(y)\;\propto\;\exp\!\big((1-\alpha)A(y)\big)\cdot\operatorname{sg}\!\Big(\pi_{k}(y)^{-(1-\alpha)}\Big).$ (7)
+Batch normalization of escort weights.
+The escort density $\rho_{\alpha}(y)\propto\pi_{k}(y)^{\alpha}P^{*}(y)^{1-\alpha}$ involves an unknown normalizing constant $Z_{\alpha}$.
+In practice we do not compute $Z_{\alpha}$ explicitly; instead, we apply batch-wise zero-centering and $z$-scoring of $\omega_{\alpha}$ over each mini-batch.
+This is equivalent to estimating the normalizer locally within the sample batch, and ensures that the estimated weights satisfy the mean constraint $\mathbb{E}[\omega_{\alpha}]=1$ at the sample level.
+The subsequent mean subtraction in the OPO update (Remark 6.2) absorbs any residual bias.
+4.2 Intrinsic Dissipation Energy as Quadratic Fluctuation Cost
+We stabilize the update by penalizing the fluctuation amplitude $v=\pi/\pi_{k}-1$ in $L^{2}(\pi_{k})$:
+$\mathcal{E}_{\mathrm{diss}}(v):=\frac{\mu}{2}\|v\|_{L^{2}(\pi_{k})}^{2}=\frac{\mu}{2}\mathbb{E}_{\pi_{k}}[v(y)^{2}].$ (8)
+This is exactly the Pearson $\chi^{2}$ geometry in ratio coordinates: $\mathbb{E}_{\pi_{k}}[v^{2}]=\mathbb{E}_{\pi_{k}}\!\left[(\pi/\pi_{k}-1)^{2}\right]$.
+Physical motivation.
+Near the anchor, a quadratic fluctuation cost matches the universal Gaussian-fluctuation principle from near-equilibrium thermodynamics: small departures from equilibrium incur a quadratic free-energy cost [19].
+4.3 OPO as a Constrained Work–Dissipation Balance
+Define the work–dissipation functional
+$v^{*}=\operatorname*{arg\,max}_{v\in L^{2}(\pi_{k})}\left\{\underbrace{\langle\omega_{\alpha},v\rangle}_{\text{external work}}\;-\;\underbrace{\frac{\mu}{2}\|v\|^{2}}_{\text{intrinsic dissipation}}\right\}\quad\text{s.t.}\quad\mathbb{E}_{\pi_{k}}[v]=0.$ (9)
+Equivalently, in loss-minimization form,
+$\boxed{\mathcal{L}_{\mathrm{OPO}}(v)=-\mathbb{E}_{\pi_{k}}[\omega_{\alpha}(y)\,v(y)]+\frac{\mu}{2}\mathbb{E}_{\pi_{k}}[v(y)^{2}],\qquad\mathbb{E}_{\pi_{k}}[v]=0.}$ (10)
+Orthogonality of the two geometries.
+$\alpha$ enters only through $\omega_{\alpha}$ (first-order drive); $\mu$ enters only through the quadratic curvature (second-order geometry). Sampling and optimization are structurally orthogonal.
+4.4 Closed-Form Solution (Linear Response Law)
+Theorem 4.3 (Closed-Form Update).
+The unique optimizer of Eq. (9) is
+$\boxed{v^{*}(y)=\frac{1}{\mu}\Big(\omega_{\alpha}(y)-1\Big).}$ (11)
 Proof.
-Direct differentiation yields a constant Hessian $\mu I$, which is independent of $\alpha$. ∎
-Remark 5.3 (Scope of Orthogonality).
-The structural orthogonality in Theorem 5.2 holds exactly in the ratio function space $L^{2}(\pi_{\mathrm{ref}})$.
-In parameter space $\theta$, interaction may arise through policy parametrization, since $\nabla_{\theta}v=\nabla_{\theta}(\exp(\Delta)-1)$ depends on the training trajectory influenced by $\alpha$.
-However, this coupling is a property of the parametrization, not the intrinsic objective curvature: the Hessian $\nabla^{2}_{v}\mathcal{L}=\mu I$ remains structurally independent of $\alpha$ regardless of the parameter mapping.
-A natural concern is whether the stop-gradient operator in $\omega_{\alpha}$ (Remark 4.3) introduces bias. The following lemma shows it does not, at the anchor point:
-Lemma 5.4 (Exactness of Surrogate Gradient at Anchor Point).
-Let $\mathcal{L}_{\mathrm{true}}(\theta)$ be the exact objective without stop-gradient on $\omega_{\alpha}$, and $\mathcal{L}_{\mathrm{OPO}}(\theta)$ the surrogate with $\operatorname{sg}$.
-Under on-policy anchoring $\pi_{\mathrm{ref}}=\pi_{\mathrm{old}}$, at the start of each iteration ($\theta=\theta_{\mathrm{old}}$):
-$\left.\nabla_{\theta}\mathcal{L}_{\mathrm{OPO}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}=\left.\nabla_{\theta}\mathcal{L}_{\mathrm{true}}(\theta)\right|_{\theta=\theta_{\mathrm{old}}}$ (17)
-Furthermore, the gradient error grows only as $\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$.
-Proof.
-By the product rule, the true function-space gradient contains a cross-term $(\nabla_{v}\omega_{\alpha})\,v$ that the surrogate omits.
-At the anchor point $\pi_{\theta}=\pi_{\mathrm{old}}$, the ratio offset is exactly $v=\pi_{\theta}/\pi_{\mathrm{old}}-1=0$, so this cross-term vanishes.
-For $\theta\neq\theta_{\mathrm{old}}$, the error is bounded by $\|(\nabla_{v}\omega_{\alpha})\,v\|=\mathcal{O}(\|v\|)=\mathcal{O}(\|\pi_{\theta}-\pi_{\mathrm{old}}\|)$, which is small within the trust region.
+Introduce a Lagrange multiplier $\lambda$ for $\mathbb{E}_{\pi_{k}}[v]=0$.
+Stationarity gives $0=-\omega_{\alpha}(y)+\mu v(y)+\lambda$.
+Taking $\mathbb{E}_{\pi_{k}}[\cdot]$ and using $\mathbb{E}_{\pi_{k}}[v]=0$ yields $\lambda=\mathbb{E}_{\pi_{k}}[\omega_{\alpha}]=1$.
+Substituting back gives $v^{*}=(\omega_{\alpha}-1)/\mu$.
 ∎
-This confirms that stop-gradient is not an engineering hack but a first-order exact approximation: the surrogate and true objectives agree to first order at every on-policy anchor.
-Under sufficiently small trust-region updates (i.e., $\|v\|$ bounded by the $\chi^{2}$ penalty), the cumulative bias across iterations remains second-order small, since each anchor reset re-zeroes the first-order error term.
-Corollary 5.5 (Global Contraction and Linear Convergence).
-Under the assumption that the policy class can represent $v^{*}$ and that local linearization of the map $\theta\mapsto v_{\theta}$ holds near the optimum, gradient descent in $v$-space on the OPO objective follows a linear dynamical system:
-$v_{k+1}-v^{*}=(1-\eta\mu)(v_{k}-v^{*})$ (18)
-where $v^{*}=\omega_{\alpha}/\mu$.
-For step size $0<\eta<2/\mu$, this system exhibits global contraction to the unique equilibrium.
-The contraction rate $|1-\eta\mu|$ depends solely on the optimization stiffness $\mu$, confirming that convergence speed is structurally decoupled from the target distribution sharpness $\alpha$.
-We use the term “orthogonalized” to emphasize this control decoupling.
-Remark 5.6 (Contrast with KL-Based Methods).
-In KL-regularized objectives (e.g., DPO), the effective loss locally resembles a logistic function $\mathcal{L}\approx-\log\sigma(\text{margin})$.
-The local curvature scales as $\beta^{2}\sigma(m)(1-\sigma(m))$, where $m$ is the logit margin.
-This curvature depends on both the temperature $\beta$ and the current data margin $m$.
-Thus, changing $\beta$ simultaneously alters both the gradient magnitude and the local stability profile, entangling the two design axes.
-OPO’s constant curvature $\mu$ avoids this data-dependent instability.
-5.2 Comparison with Parametric L2 Regularization
-A natural question is whether OPO differs fundamentally from standard Policy Gradient with L2 parameter regularization (weight decay).
-$\mathcal{L}_{\text{L2-PG}}=-\mathbb{E}[A\log\pi_{\theta}]+\frac{\lambda}{2}\|\theta-\theta_{0}\|^{2}$ (19)
-While both methods induce a quadratic penalty, they operate in fundamentally different spaces:
-  1. 1.
-    Functional vs.
-Parametric Prior: OPO’s penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ acts in the function space of probability ratios.
-It penalizes deviations only where the policy assigns probability mass.
-L2 parameter regularization acts in the weight space.
-Two different policies can have the same L2 distance, and small weight changes can cause large output deviations (and vice-versa).
-  2. 2.
-    Adaptive Constraint: OPO’s regularization provides a “linear restoring force” directly on the probability ratio.
-If the model is confident (large $v$), the $\chi^{2}$ penalty generates a strong, specific gradient signal to pull that specific output back.
-Parametric L2 is blind to the output; it uniformly decays all weights regardless of their contribution to the current output confidence.
-  3. 3.
-    Learning Dynamics: Empirically, Functional $\chi^{2}$ regularization allows the model to shift mass aggressively where advantages are high, while maintaining stability.
-Weight decay suppresses all weight magnitudes, which can inadvertently hinder the learning of sharp, high-confidence reasoning chains required for math tasks.
-5.3 Quantitative Gradient Analysis
-We provide a quantitative comparison of the gradient signal in high-confidence regimes, proving the non-saturation property.
-Gradient Saturation in DPO.
-For logistic-based losses (DPO), the gradient magnitude with respect to the margin $m$ is $|\nabla_{m}\mathcal{L}|=|\beta\sigma(m)(1-\sigma(m))|$. Since $\sigma(1-\sigma)\leq 1/4$:
-$|\nabla_{\text{DPO}}|\leq\frac{\beta}{4}$ (20)
-Crucially, as the margin $m\to\infty$, $|\nabla_{\text{DPO}}|\to 0$ exponentially.
-Linear Gradient Dynamics in OPO.
-For OPO, the gradient is $\nabla_{v}\mathcal{L}=\mu v-\omega_{\alpha}$.
-Let $v^{*}=\omega_{\alpha}/\mu$.
-The gradient magnitude is exactly proportional to the distance from equilibrium: $|\nabla_{v}\mathcal{L}|=\mu|v-v^{*}|$.
-Therefore, for any non-equilibrium state ($|v-v^{*}|\geq\delta$):
-$|\nabla_{\text{OPO}}|\geq\mu\delta$ (21)
-This guarantees a linear driving force that does not vanish arbitrarily.
-The practical consequence is stark: in mathematical reasoning, models frequently assign near-1.0 probability to correct intermediate steps.
-KL-based methods penalize these high-confidence correct outputs by exponentially shrinking the learning signal, whereas OPO maintains a constant reward-driven pull via its linear dynamics—allowing the model to continue refining even its most confident predictions.
+Policy update.
+Since $\pi=\pi_{k}(1+v)$,
+$\frac{\pi_{k+1}(y)}{\pi_{k}(y)}=1+\frac{1}{\mu}\Big(\omega_{\alpha}(y)-1\Big).$ (12)
+Non-saturating linear dynamics.
+The function-space gradient is $\nabla_{v}\mathcal{L}_{\mathrm{OPO}}=-\omega_{\alpha}+\mu v$, which is linear in $v$ with constant curvature $\nabla_{v}^{2}\mathcal{L}=\mu I$.
+Unlike sigmoid/logistic objectives, this does not exhibit exponential saturation.
+Figure 1: Work–dissipation physics picture of OPO.
+The anchor policy $\pi_{k}$ sits at the minimum of a quadratic fluctuation energy ($\chi^{2}$ geometry).
+The escort field $\rho_{\alpha}$ induces the driving weight $\omega_{\alpha}=\rho_{\alpha}/\pi_{k}$.
+The perturbed policy $\pi$ corresponds to a ratio fluctuation $v=\pi/\pi_{k}-1$.
+The conservation law $\mathbb{E}_{\pi_{k}}[v]=0$ forces an orthogonal projection, yielding the centered response $v^{*}=(\omega_{\alpha}-1)/\mu$.
+In statistical mechanics [18], this is the fluctuation–dissipation theorem: response equals susceptibility ($1/\mu$) times centered driving force.
+5 Mirror Descent Interpretation
+We show that the work–dissipation principle is equivalently a mirror descent step with a Euclidean mirror map.
+5.1 Euclidean Mirror Map and $\chi^{2}$ Bregman Divergence
+Define the mirror map $\Psi(v):=\frac{1}{2}\|v\|_{L^{2}(\pi_{k})}^{2}$. The induced Bregman divergence is
+$D_{\Psi}(\pi\|\pi_{k})=\frac{1}{2}\mathbb{E}_{\pi_{k}}\!\left[\left(\frac{\pi}{\pi_{k}}-1\right)^{2}\right],$ (13)
+which is the Pearson $\chi^{2}$ divergence in ratio coordinates.
+5.2 Orthogonal Mirror Step
+A mirror/proximal step driven by $\omega_{\alpha}$ is
+$v^{*}=\operatorname*{arg\,min}_{v\in\mathcal{H}_{0}}\left\{-\langle\omega_{\alpha},v\rangle+\frac{\mu}{2}\|v\|^{2}\right\},$ (14)
+which is identical to the work–dissipation formulation (9).
+Thus mirror descent and the free-energy principle define the same update.
+$\alpha$ affects only the drive $\omega_{\alpha}$; $\mu$ affects only the proximal curvature.
+6 Hilbert Projection Interpretation
+We re-derive the same objective from the Hilbert space projection theorem, requiring no mirror maps.
+6.1 Hilbert Space and Metric Modulation
+Let $\mathcal{H}=L^{2}(\pi_{k})$ with inner product $\langle f,g\rangle=\mathbb{E}_{\pi_{k}}[fg]$. Define the multiplication operator (diagonal preconditioner)
+$(\mathcal{M}_{\alpha}f)(y):=\omega_{\alpha}(y)\,f(y).$ (15)
+Lemma 6.1 (Boundedness, Self-Adjointness, Positivity).
+If $\omega_{\alpha}\in L^{\infty}(\pi_{k})$ and $\omega_{\alpha}>0$ a.s., then $\mathcal{M}_{\alpha}$ is bounded, self-adjoint, and positive.
+The induced inner product $\langle f,g\rangle_{\alpha}:=\mathbb{E}_{\rho_{\alpha}}[fg]$ is the escort-modulated geometry.
+Thus $\alpha$-escort is a smooth metric modulation: it preserves support and reweights directions in function space through a new local inner product, in contrast to hard masking which truncates support.
+6.2 Projection Theorem Gives the Same Closed Form
+Rewrite Eq. (9) by completing the square:
+$\langle\omega_{\alpha},v\rangle-\frac{\mu}{2}\|v\|^{2}=-\frac{\mu}{2}\left\|v-\frac{1}{\mu}\omega_{\alpha}\right\|^{2}+\frac{1}{2\mu}\|\omega_{\alpha}\|^{2}.$ (16)
+Hence
+$v^{*}=\operatorname*{arg\,min}_{v\in\mathcal{H}_{0}}\left\|v-\frac{1}{\mu}\omega_{\alpha}\right\|^{2}.$ (17)
+By the Hilbert projection theorem,
+$v^{*}=\frac{1}{\mu}\mathcal{P}_{\mathcal{H}_{0}}(\omega_{\alpha})=\frac{1}{\mu}\big(\omega_{\alpha}-\langle\omega_{\alpha},1\rangle\,1\big)=\frac{1}{\mu}(\omega_{\alpha}-1),$ (18)
+recovering Theorem 4.3. The subtraction of $1$ is the unique orthogonal correction enforced by conservation.
+Remark 6.2 ($z$-Score Normalization as Conservation-Law Projection).
+The centering $\omega_{\alpha}\mapsto\omega_{\alpha}-\mathbb{E}[\omega_{\alpha}]$ is exactly $\mathcal{P}_{\mathcal{H}_{0}}$—the mean-subtraction step of $z$-score normalization.
+When $\omega_{\alpha}(y)=A(y)$ ($\alpha\to 1$), the solution $v^{*}=(A-\mathbb{E}[A])/\mu$ is the $z$-scored advantage scaled by susceptibility.
+Thus, advantage $z$-scoring is not a variance-reduction heuristic: it is the unique orthogonal projection enforcing probability conservation in $L^{2}(\pi_{k})$.
+Remark 6.3 (Contrast with Fisher–Rao (KL) Geometry).
+In KL-based methods, the natural metric is the Fisher–Rao metric $g_{ij}^{\mathrm{FR}}=\mathbb{E}_{\pi}[\partial_{i}\log\pi\cdot\partial_{j}\log\pi]$, whose curvature depends on $\pi$.
+In contrast, $\chi^{2}$ geometry operates in $L^{2}(\pi_{k})$ with constant inner product:
+KL / Fisher–Rao $\chi^{2}$ / $L^{2}(\pi_{k})$
+Inner product Fisher metric (policy-dependent) $L^{2}(\pi_{k})$ (fixed at anchor)
+Hessian $\nabla^{2}\propto\beta^{2}\sigma(1{-}\sigma)$ $\nabla^{2}=\mu I$ (constant)
+Perturbation theory Nonlinear (exponential) Linear (first-order exact)
+7 Implementable Surrogate and Algorithm
+7.1 Log-Ratio Surrogate for LLMs
+LLMs output log-probabilities. Let $\Delta_{\theta}(y):=\log\pi_{\theta}(y)-\log\pi_{k}(y)$. Since $v=e^{\Delta}-1\approx\Delta$ for small trust regions, the practical surrogate is
+$\boxed{\mathcal{L}_{\mathrm{OPO}}^{\log}(\theta)=-\mathbb{E}_{\pi_{k}}\big[\omega_{\alpha}(y)\,\Delta_{\theta}(y)\big]+\frac{\mu}{2}\mathbb{E}_{\pi_{k}}\big[\Delta_{\theta}(y)^{2}\big].}$ (19)
+Anchoring $\pi_{k}=\pi_{\mathrm{old}}$ resets $\Delta$ near zero each iteration, supporting the local approximation.
 7.2 Algorithm and Flowchart
-We provide a visual overview of the OPO process in Figure 1 and the detailed procedure in Algorithm 1.
 Current Policy $\pi_{\theta}$ Sampling $(x,y)$ Evaluation $A(y)$ OPO Decomposition 1.
@@
 Penalty $\mu\,v^{2}$ Update $\theta$ Next Iteration
-Figure 1: Flowchart of the Orthogonalized Policy Optimization (OPO) framework.
-The “OPO Decomposition” block (green) embodies the two-geometry structure: the $\alpha$-divergence projection (step 2) governs sampling, while the $\chi^{2}$ penalty (step 3) governs optimization.
-These two axes are structurally orthogonal.
+Figure 2: Flowchart of the OPO framework.
+The “OPO Decomposition” block (green) embodies the work–dissipation structure: the $\alpha$-escort (step 2) shapes the external work, while the $\chi^{2}$ penalty (step 3) shapes the intrinsic dissipation.
+The two axes are structurally orthogonal.
 Algorithm 1 Orthogonalized Policy Optimization (OPO)
@@
 4:  Rollout: Sample $(x,y)\sim\pi_{\theta}$; compute rewards $R(x,y)$
-5:  Advantage: $A(y)\leftarrow\mathrm{Normalize}(R(x,y))$ $\triangleright$ Group-normalized
+5:  Advantage: $A(y)\leftarrow\mathrm{zscore}(R(x,y))$ $\triangleright$ Group $z$-score (Remark 6.2)
 6:  Escort Weight: $\omega_{\alpha}(y)\leftarrow\exp\!\big((1{-}\alpha)\,A(y)\big)\cdot\operatorname{sg}\!\big(\pi_{\mathrm{ref}}(y)^{-(1-\alpha)}\big)$ $\triangleright$ Eq. 7
 7:  Log-Ratio: $\Delta_{\theta}(y)\leftarrow\log\pi_{\theta}(y)-\log\pi_{\mathrm{ref}}(y)$
-8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$ $\triangleright$ $y$: complete response sequence
+8:  Loss: $\mathcal{L}\leftarrow\frac{1}{N}\sum_{y}\big[-\omega_{\alpha}(y)\,\Delta_{\theta}(y)+\frac{\mu}{2}\,\Delta_{\theta}(y)^{2}\big]$
 9:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
 10:end for
+8 Theoretical Analysis
+8.1 Constant Curvature and Orthogonal Control
+Theorem 8.1 (Decoupling of Sampling and Optimization Geometry).
+Consider the OPO objective in $v$-space. Then:
+  1. 1.
+    The first-order drive depends on $\alpha$ only through $\omega_{\alpha}$: $\nabla_{v}\mathcal{L}_{\mathrm{OPO}}=-\omega_{\alpha}+\mu v$.
+  2. 2.
+    The second-order curvature is constant and independent of $\alpha$: $\nabla_{v}^{2}\mathcal{L}_{\mathrm{OPO}}=\mu I$.
+Changing $\alpha$ reshapes the target (what to learn) without altering curvature (how to learn). Changing $\mu$ alters curvature without reshaping the target.
+8.2 Global Contraction in $v$-Space
+Gradient descent $v^{(m+1)}=v^{(m)}-\eta(\mu v^{(m)}-\omega_{\alpha})$ satisfies
+$v^{(m+1)}-v^{*}=(1-\eta\mu)\,(v^{(m)}-v^{*}).$ (20)
+Corollary 8.2 (Linear Rate).
+For $0<\eta<2/\mu$, $\|v^{(m)}-v^{*}\|\leq|1-\eta\mu|^{m}\,\|v^{(0)}-v^{*}\|$. The contraction rate depends only on $\mu$ and is independent of $\alpha$.
+8.3 Why KL-Based Objectives Saturate
+In KL-regularized preference optimization, local curvature scales like $\sigma(m)(1-\sigma(m))$ and vanishes exponentially as $|m|\to\infty$.
+By contrast, OPO has constant curvature $\mu I$ and linear restoring force $\|\nabla_{v}\mathcal{L}\|=\|\mu(v-v^{*})\|$, which remains proportional to distance-to-equilibrium.
+Figure 3 visualizes this contrast on a binary logit toy model ($z_{\mathrm{ref}}=0$).
+Panel (a) plots $|\partial\mathcal{L}/\partial z|$ versus the model confidence $p=\sigma(z)$: SFT, DPO, GRPO, DAPO, and GSPO all decay to zero in the high-confidence regime, while OPO’s $\chi^{2}$ gradient grows linearly.
+Panel (b) zooms into $p\geq 0.5$ on a logarithmic scale, revealing that clipped objectives (GRPO, DAPO, GSPO) suffer a hard cliff to machine-epsilon once the ratio exits the clip window, and DPO saturates exponentially.
+OPO is the only objective whose gradient magnitude remains $O(1)$ throughout.
+Figure 3: Toy gradient saturation experiment (binary logit, $z_{\mathrm{ref}}=0$).
+(a) Gradient magnitude vs.
+model confidence: OPO ($\chi^{2}$, green) exhibits linear growth while all other objectives saturate.
+(b) Log-scale zoom on $p\geq 0.5$: clipped objectives cliff-drop by up to 9 orders of magnitude; OPO maintains non-vanishing gradients throughout.
+8.4 Trust-Region Duality
+Proposition 8.3 ($\chi^{2}$ Trust-Region Duality).
+The constrained problem $\max_{v\in\mathcal{H}_{0}}\langle\omega_{\alpha},v\rangle$ s.t.
+$\|v\|^{2}\leq\varepsilon$ has Lagrangian relaxation $\max_{v\in\mathcal{H}_{0}}\langle\omega_{\alpha},v\rangle-\frac{\mu}{2}\|v\|^{2}$, with $\mu$ as the multiplier.
+Moreover, $\|v\|^{2}=\mathbb{E}_{\pi_{k}}[(\pi/\pi_{k}-1)^{2}]$ is the Pearson $\chi^{2}$ divergence.
+8.5 Positivity Safeguards
+The closed-form linear response $v^{*}(y)=(\omega_{\alpha}(y)-1)/\mu$ is derived in the tangent space of the probability manifold at $\pi_{k}$, and only guarantees first-order stationarity of the proximal objective.
+It does not enforce $\pi_{k}(y)(1+v^{*}(y))\geq 0$ outside a local neighborhood.
+When $\omega_{\alpha}(y)\ll 1$ and $\mu$ is small, one may have $v^{*}(y)<-1$, violating positivity.
+Proposition 8.4 (Local Positivity Condition).
+The update $\pi_{k+1}=\pi_{k}(1+v^{*})$ is a valid probability distribution whenever $\mu\geq\|\omega_{\alpha}\|_{\infty}-1$, since then $v^{*}(y)=(\omega_{\alpha}(y)-1)/\mu\geq-1$ for all $y$.
+In practice, positivity is maintained by one or more of: (i) choosing $\mu$ large enough relative to the escort weight range; (ii) clipping $v^{*}$ to $1+\epsilon,\,+\infty)$; (iii) projecting onto the probability simplex after the update.
+This is consistent with standard trust-region Taylor approximations in policy optimization and does not alter local convergence properties (Corollary 8.2).
 9 Experiments
@@
   • GRPO (Group Relative Policy Optimization) [4]: Standard token-level policy gradient with group-normalized advantages. Uses vanilla PPO loss mode without a value function critic.
-  • GSPO (Group Sentence-level Policy Optimization): A variant of GRPO that applies advantage normalization at the sentence/step level rather than the token level, aiming for more granular credit assignment.
-  • DAPO (Dynamic Advantage Policy Optimization): An advanced baseline incorporating four key mechanisms:
+  • GSPO (Group Sequence Policy Optimization) [21]: Applies sequence-level importance ratios rather than token-level ratios, aiming for more granular credit assignment at the trajectory level.
+  • DAPO (Decoupled Clip and Dynamic sAmpling Policy Optimization) [20]: An advanced baseline incorporating four key mechanisms:
     1. 1.
@@
     2. 2.
-      Dynamic Sampling: dynamically filters "all-correct" or "all-wrong" samples that contribute zero advantage, focusing training on informative samples.
+      Dynamic Sampling: dynamically filters “all-correct” or “all-wrong” samples that contribute zero advantage, focusing training on informative samples.
     3. 3.
@@
 Figure 5: Gradient norm comparison. OPO maintains healthy gradient norms (0.9) throughout training, while DAPO (0.22) and GSPO (0.50) exhibit gradient decay.
+Gradient Saturation Analysis.
+Figure 6 provides direct evidence for OPO’s non-saturating property.
+Panel (a) plots gradient norms against mean reward (a confidence proxy).
+As the model becomes more confident (higher reward), GRPO and GSPO gradients exhibit a decaying trend, and DAPO gradients remain near zero.
+In contrast, OPO maintains elevated gradient norms across all confidence levels.
+Panel (b) addresses the fact that different methods operate at different advantage scales (e.g., GRPO $z$-scores advantages to $\sigma\!\approx\!2$, whereas OPO uses raw advantages with $\sigma\!\approx\!0.2$).
+To enable a fair comparison, we plot the gradient response efficiency $\|\nabla\|/\bigl(\tfrac{1}{2}\text{AdvRange}\bigr)$ over training steps, normalizing gradient magnitude by the advantage signal amplitude.
+OPO sustains an efficiency of ${\sim}1.0$$1.5$ throughout training, approximately $5{\times}$ higher than GRPO (${\sim}0.2$$0.3$) and DAPO (${\sim}0.1$).
+This confirms the linear response prediction of Theorem 4.3: under $\chi^{2}$ geometry, each unit of advantage signal produces a proportional gradient update, whereas KL-based objectives attenuate the signal through their policy-dependent curvature.
+Figure 6: Gradient saturation analysis.
+(a) Gradient norm vs.
+mean reward (confidence proxy): OPO maintains non-saturating gradients as confidence grows, while baselines decay.
+(b) Gradient response efficiency (gradient norm normalized by advantage amplitude) over training: OPO sustains ${\sim}5{\times}$ higher gradient-per-signal than baselines, confirming the constant-curvature prediction of $\chi^{2}$ geometry.
 Entropy Dynamics.
@@
 The experimental results confirm that OPO’s orthogonalized design—combining $\alpha$-weighted sampling with $\chi^{2}$ optimization geometry—prevents gradient saturation and enables continued learning in high-confidence regimes.
+The gradient saturation analysis (Figure 6) provides the most direct evidence: OPO’s gradient magnitude remains proportional to advantage amplitude (linear response), while KL-based baselines exhibit suppressed gradients regardless of the driving signal.
 10 Discussion
@@
   • OPO: moderate $\alpha$ (peak-seeking escort) combined with Euclidean ($\chi^{2}$) geometry—non-saturating gradients with advantage-aware sampling.
+Geometry Comparison.
+Table 2 summarizes the structural differences among representative methods.
+Table 2: Comparison of sampling and optimization geometries across alignment methods.
+Method Sampling Geometry Optimization Geometry
+PPO / GRPO implicit KL KL curvature (ratio clip)
+DPO KL KL curvature (sigmoid)
+OPO (Ours) $\alpha$-escort (tunable) $\chi^{2}$ curvature (constant $\mu I$)
+Geometric Interpretation of Entropy Preservation.
+From a geometric perspective, KL geometry inherits a logarithmic penalty on low-probability outcomes: $\log\pi(y)$ diverges as $\pi(y)\to 0$, strongly suppressing entropy by penalizing any residual mass on unlikely tokens.
+In contrast, $\chi^{2}$ geometry penalizes deviations quadratically in ratio space: the cost $(\pi/\pi_{k}-1)^{2}$ remains finite even when $\pi(y)\ll\pi_{k}(y)$, which attenuates the shrinkage of low-probability mass and thus helps maintain entropy.
+This geometric explanation complements the sequence-level vs.
+token-level distinction discussed in the experiments.
 Computational Cost.
@@
   • For on-policy training where samples come from $\pi_{\mathrm{ref}}=\pi_{\text{old}}$, the batch itself provides an unbiased estimate.
+  • Mini-batch anchoring.
+Within one outer iteration, $\pi_{\mathrm{ref}}$ is frozen at the policy snapshot taken before the first mini-batch gradient step.
+All subsequent mini-batches in the same iteration share this anchor, ensuring that the log-ratio $\Delta_{\theta}=\log\pi_{\theta}-\log\pi_{\mathrm{ref}}$ is evaluated against a consistent reference.
+This is analogous to PPO’s “epoch” convention and guarantees that $\Delta_{\theta}$ starts near zero at the beginning of each outer iteration, supporting the local linear approximation $v\approx\Delta$.
   • For off-policy settings, importance sampling or a separate reference sample set can be used.
@@
 11 Conclusion
-We have presented Orthogonalized Policy Optimization (OPO), a framework grounded in two independent geometries: an $\alpha$-divergence escort projection that shapes sampling, and a Euclidean mirror map (Pearson $\chi^{2}$) that shapes optimization.
-The same loss emerges from two complementary derivations—mirror descent and ratio-space regression—confirming it is uniquely determined by the geometry choice.
-The resulting objective has a constant Hessian, linear non-saturating gradients, and global contraction, addressing a key limitation of KL-based methods in high-confidence reasoning tasks.
-We believe OPO provides a principled foundation for future work on robust and scalable LLM alignment.
+We have presented Orthogonalized Policy Optimization (OPO), grounded in a unified work–dissipation principle: the policy update maximizes external work from an $\alpha$-escort sampling field while paying a quadratic dissipation cost in $\chi^{2}$ ratio geometry.
+This single variational principle admits three equivalent interpretations—mirror descent, Hilbert-space projection, and linear response—confirming that OPO is the canonical quadratic proximal response in ratio geometry.
+The resulting framework has constant Hessian, linear non-saturating gradients, global contraction, and reveals advantage $z$-scoring as a conservation-law projection.
+Experiments on mathematical reasoning tasks validate that OPO outperforms KL-based methods where gradient saturation limits further improvement.
 Appendix A Theoretical Proofs and Derivations
@@
 Let $\delta=\|\Delta_{\theta}\|_{\infty}$.
-For sufficiently small $\delta$ (in particular $\delta<1$, which is enforced by the $\chi^{2}$ trust-region penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$ bounding $\|\Delta\|_{\infty}$):
+For sufficiently small $\delta$ (in particular $\delta<1$, which holds under the small trust-region regime maintained by on-policy anchoring and the $\chi^{2}$ penalty $\frac{\mu}{2}\mathbb{E}[v^{2}]$):
 $|v_{\theta}-\Delta_{\theta}|\leq\frac{1}{2}\delta^{2}e^{\delta},\quad|v_{\theta}^{2}-\Delta_{\theta}^{2}|\leq\delta^{3}e^{2\delta}$ (21)